### v2

# From Prejudice to Parity: A New Approach to Debiasing Large Language Model Word Embeddings

Abstract
Embeddings play a pivotal role in the efficacy of Large Language Models. They are the bedrock on which these models grasp contextual relationships and foster a more nuanced understanding of language and consequently perform remarkably on a plethora of complex tasks that require a fundamental understanding of human language. Given that these embeddings themselves often reflect or exhibit bias, it stands to reason that these models may also inadvertently learn this bias. In this work, we build on the seminal previous work and propose DeepSoftDebias, an algorithm that uses a neural network to perform ‘soft debiasing’. We exhaustively evaluate this algorithm across a variety of SOTA datasets, accuracy metrics, and challenging NLP tasks. We find that DeepSoftDebias outperforms the current state-of-the-art methods at reducing bias across gender, race, and religion.
1 Introduction
Word embeddings are a foundational element in the architecture of large language models (LLMs). They act as the basis for these models to understand and subsequently, generate human-like language. However, it has been shown that these models rely on word embeddings, which themselves may reflect or exhibit bias Papakyriakopoulos et al. (2020); Dev et al. (2020). Given the exponential increase in the use of large language models on a plethora of downstream tasks, these representations can amplify bias and result in discriminatory actions, especially when it comes to the fields of education, healthcare, and justice. Existing work in this field has looked most commonly into gender bias Kotek et al. (2023); Bordia and Bowman (2019); de Vassimon Manela et al. (2021), racial bias Mozafari et al. (2020); Omiye et al. (2023); Tang et al. , and religious bias Baligudam (2022); Kirk et al. (2021). In this work, we build on the seminal work of Gonen and Goldberg (2019), which brought attention to the inherent biases present in traditional GloVe embeddings Pennington et al. (2014). This study prompted the NLP community to reevaluate the fundamental choices underlying our word representation models. Specifically, we present DeepSoftBias: an algorithm that furthers the application of their methodology, by diverging from the conventional GloVe embeddings and delving into the word embeddings produced by the best-performing models on the Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2022) leaderboard. By employing these advanced embeddings on the same set of words as used in GloVe embeddings, we seek to investigate whether these state-of-the-art (SoTA) models inherently exhibit reduced bias. Our primary objective is twofold: first, to de-bias the embeddings from these selected models, and second, to rigorously assess the effectiveness of the bias removal process. Our proposed approach, DeepSoftDebias, is an innovative methodology to de-bias LLM word embeddings which involves integrating a neural network into the soft debiasing approach developed by Bolukbasi et al. (2016). This novel amalgamation is driven by the aspiration to enhance the debiasing process and contribute to the ongoing discourse on creating fair and ethically sound language models. To this end, our work answers the following research questions:
  1. RQ1: Compared to traditional methods, does our proposed methodology attain better performance metrics when it comes to debiasing SOTA model embeddings?
  2. RQ2: How do parameters of the model (size, complexity) interact with various SOTA debiasing techniques? What effect do they have on each other?
  3. RQ3: To what extent do various SOTA debiasing techniques influence the performance of models on different downstream tasks?
  4. RQ4: How does the type of bias (gender, race, religion) affect the effectiveness of the debiasing process?
To answer the above questions, we make the following contributions through this research:
Our Contributions
[leftmargin=1mm] We provide, to the best of our knowledge, the first comprehensive study of how various debiasing methods work on SoTA LLM word embeddings We present a novel methodology, DeepSoftDebias, for debiasing LLM word embeddings, which beats SoTA debiasing methods across multiple bias formats including gender, race, and religion. We perform an exhaustive quantitative analysis, establishing SoTA baselines and leveraging multiple evaluation metrics to provide a comparison against accessible SoTA baselines
We illustrate our pipeline in Fig. 1. We find that DeepSoftDebias not only outperforms the state-of-the-art methods at reducing bias across gender, race, and religion but also does so while preserving the full information of the original embedding (which is an additional improvement on previous methods). Further, we find that model performance on challenging downstream tasks like NER and sentiment analysis remains largely unaffected when we test using our debiased embeddings.
Figure 1: A step-by-step visualization of the pipeline for DeepSoftDebias. Our pipeline has 3 major components, Initial Word Vector Generation, Debiasing, and Quantitative Analysis. The Debiasing stage leverages the DeepSoftDebias network.
2 Data
This study leverages several datasets to examine and address biases in word embeddings and language models, focusing on the representation and perpetuation of stereotypes within these systems.
2.1 L2-Reddit Corpus
We utilize the L2-Reddit [^1] https://github.com/ellarabi/reddit-l2 Rabinovich et al. (2018) corpus, a collection of Reddit posts and comments by both native and non-native English speakers, featuring approximately 56 million sentences from the United States. This dataset serves as the foundation for training the word embeddings, aiming to capture the nuanced and inherently biased linguistic patterns present in social media discourse. The corpus provides a rich linguistic resource for analyzing and understanding the initial biased word embeddings. In our study, we employ the Reddit L2 corpus as the source for our initial Word2Vec Mikolov et al. (2013) word embeddings. Subsequently, we leverage the vocabulary derived from these word vectors to obtain the word embeddings from the Language Model Models (LLMs).
2.2 StereoSet
StereoSet Nadeem et al. (2020) stands out as a critical dataset for measuring stereotype bias in language models, containing around 17,000 sentences across demographic dimensions like gender, race, religion, and profession. It introduces the Context Association Tests (CAT) for evaluating model preferences and biases, providing a structured approach to assess and quantify biases in popular models like BERT Devlin et al. (2019), GPT-2 Radford et al. (2019), RoBERTa Liu et al. (2019), and XLNet Yang et al. (2020). The dataset’s creation involved a rigorous crowdsourcing process, ensuring a wide coverage of stereotypes and anti-stereotypes to challenge and evaluate the fairness of NLP systems. In our words we use the stereoset dataset to benchmark our debiasing method against all the other debiasing methods. To accomplish this, we compute the average sentence vector for all sentences and their respective contexts. We then compare these vectors against each other to derive the stereotype score.
2.3 CrowS-Pairs
CrowS-Pairs Nangia et al. (2020), designed to assess social biases in masked language models (MLMs), comprises 1,508 examples covering nine bias types, including race, religion, and age. It contrasts sentences related to historically disadvantaged and advantaged groups in the U.S., with annotations from crowd workers highlighting the degree of stereotyping. Despite its noise and reliability issues, CrowS-Pairs remains a valuable resource for evaluating the bias in major language models, providing a nuanced metric that quantifies preference for stereotypical versus non-stereotypical expressions. In our study, we obtain debiased word embeddings for sentences by computing the average sentence vector for both less and more stereotypical or anti-stereotypical directions. We then compare these embeddings against each other to calculate the Crows Metric score.
3 Methodology
In this research, we delve into the domain of debiasing word embeddings, presenting both an established and a newly proposed methodology for mitigating biases in word vector representations. These biases span across gender, racial, and religious lines and are encoded inadvertently within language models.
3.1 Bias Identification and Data Structure
To quantitatively assess bias in word embeddings, we measure the projection of word vectors onto a gender-specific axis, defined by the vectorial difference between the terms ‘he’ and ‘she.’ The magnitude of this projection serves as an indicator of bias. We use a structured vocabulary with its associated vector representations from the Word2Vec model to facilitate the identification of biases. For a comprehensive evaluation, we utilize additional data files that include definitive sets of gender-associated word pairs, analogy templates that list occupational roles often linked with specific genders, and a set of neutral terms used as evaluation targets. These resources are crucial for the systematic identification and rectification of biases in word embeddings.
3.2 Soft Debiasing: The Baseline Approach
The initial method as seen in the paper Manzini et al. (2019) leverages a method called soft debiasing. We recap its algorithm in Algorithm 1. Soft debiasing involves learning a projection of the embedding matrix that preserves the inner product between biased and debiased embeddings while minimizing the projection onto the bias subspace of embeddings mentioned in 3.1 that should be neutral. Given embeddings $W$ and $N$ which are embeddings for the whole vocabulary and the subset of bias-neutral words respectively, and the bias subspace B obtained in Section 2.1, soft debiasing seeks a linear transformation A that minimizes the following objective defined in Eq. (1) as follows:
$\left\|(AW)^{T}(AW)-W^{T}W\right\|_{F}^{2}+\lambda\left\|(AN)^{T}(AB)\right\|_{F}^{2}$
Minimizing the first term preserves the inner product after the linear transformation A, and minimizing the second term minimizes the projection onto the bias subspace B of embeddings. R is a tunable parameter that balances the two objectives.
Input: Biased word embeddings ($\text{emb}_{\text{biased}}$), Bias Subspace (BiasSpace), Neutral word embeddings ($\text{emb}_{\text{neutral}}$)
Output: Debiased word embeddings
Perform Singular Value Decomposition (SVD) on $\text{emb}_{\text{biased}}$ to obtain singular values ($s$) and left singular vectors ($u$);
Precompute $t1=s\cdot u^{T}$ and $t2=u\cdot s$;
Compute norm1 as $\|t1\cdot(T^{T}\cdot T-I)\cdot t2\|_{F}$;
Compute norm2 as $\|\text{emb}_{\text{neutral}}^{T}\cdot T^{T}\cdot\text{BiasSpace}\|_{F}$;
Total loss is a weighted combination of norm1 and norm2;
Optimize transformation matrix using SGD;
Output debiased word embeddings after recomputing using $T$ and normalizing;
Algorithm 1 Transformation Matrix Approach
3.3 DeepSoftDebias: Our Proposed Approach
In the original approach introduced by Bolukbasi et al. (2016), a transformation matrix is utilized and optimized by an optimizer to enable a direct mapping between input and output embeddings. To enhance performance, we propose DeepSoftDebias. In this approach, we replace the transformation matrix with a neural network, leveraging its capability to represent a sequence of transformation matrices. This adaptation enables the algorithm to handle more complex functions mapping between input and output embeddings. We use the same loss functions as mentioned in the section 3.2. Furthermore, we transition from stochastic gradient descent (SGD Robbins and Monro (1951)) to the Adam Kingma and Ba (2017) optimizer, resulting in enhanced efficiency, speed, and optimization quality. We describe our full algorithm in Algorithm 2. While these modifications were implemented, the fundamental aspects of the method remain unaltered, ensuring minimal alterations in embeddings and preserving orthogonality with the bias space. Unlike the baseline, which relies on singular value decomposition (SVD) and incurred information loss, DeepSoftDebias preserves the full information of the original matrix. Moreover, unlike the baseline, DeepSoftDebias can handle large embedding dimensions of more than 4.5k. We demonstrate the effectiveness of DeepSoftDebias on various datasets and tasks, and show that it outperforms the state-of-the-art methods in terms of accuracy and efficiency. Fig. 2 presents a visualization of our approach to downstream testing.
Input: Biased word embeddings ($\text{emb}_{\text{biased}}$), Bias Subspace (BiasSpace), Neutral word embeddings ($\text{emb}_{\text{neutral}}$)
Output: Debiased word embeddings
Initialize neural network $NN$ with input dimension as embedding dimension and output dimension as embedding dimension;
Pass $\text{emb}_{\text{biased}}$ through $NN$ to obtain transformed embeddings;
Compute $T^{T}$ as the matrix multiplication of the transpose of outputs of $NN$ and the outputs;
Compute norm1 as $\|(T^{T}\cdot T-I)\|_{F}$;
Compute norm2 as $\|\text{emb}_{\text{neutral}}^{T}\cdot T^{T}\cdot\text{BiasSpace}\|_{F}$;
Total loss is a weighted combination of norm1 and norm2;
Optimize $NN$ using an Adam optimizer;
Output normalized embeddings obtained after passing $\text{emb}_{\text{biased}}$ through $NN$;
Algorithm 2 Neural Network Approach
Figure 2: A step-by-step visualization of our downstream testing process to effectively evaluate DeepSoftDebias
4 Effects of LLM Size and Dependency of Network Size
The debiasing performance of word embeddings depends on the size of the embeddings and the depth of the debiasing neural network, rather than the number of parameters of the language model. We observe in 11 Smaller models, such as bge-small Xiao et al. (2023) and DeBERTa-v3-base He et al. (2023) or DeBERTa-v3-large, can be debiased effectively by a single-layer neural network. Larger models, such as Llama-2 Touvron et al. (2023), Alpaca Taori et al. (2023) and Yi-6b 01.ai (2024) need a more complex debiasing neural network. For embedding dimensions around 2000, a two-layer neural network is sufficient, while for larger embedding dimensions, a three-layer neural network is required to achieve good debiasing results. Additionally, the debiasing neural network and the optimization algorithm need to be hyperparameter-tuned, such as adjusting the learning rate, to get optimal results. The hyperparameters may vary depending on the model size, the embedding dimension, and the debiasing task. Therefore, a systematic search for the best hyperparameters is necessary to ensure the effectiveness of the debiasing process.
5 Results
The results obtained are described in the following subsections.
5.1 Mean Average Cosine Similarity (MAC)
MAC Manzini et al. (2019) is a metric used to quantify semantic associations between word classes and attributes. MAC takes word embeddings, targets (representing classes), and attributes as inputs. By computing the mean cosine distance between target words and attribute sets, MAC offers a concise measure of semantic proximity. This metric provides valuable insights into the contextual semantics encoded within word embeddings. Table 1 shows that the word embeddings debiased in the direction of race and gender have comparable increases in their average MAC of  0.64, whereas word embeddings debiased in the direction of religion have an increase in MAC of around 0.61. We see that our debiasing procedure categorically moves MAC scores closer to 1.0. This indicates an increase in cosine distance. Further, the associated P-values indicate these changes are statistically significant. This demonstrates that our approach for multiclass debiasing decreases bias in the word embeddings. We provide visual representations of the efficiency of DeepSoftDebias at removing gender bias, racial bias, and religion bias in Appendix A.
Model Variant Topic BMAC NSMAC SS CMS CSS CAS
Yi Yi-6B Gender 0.148 0.964 55.372 49.620 58.970 37.250
Alpaca Alpaca-7B 0.612 0.816 53.306 48.850 57.690 37.250
BAAI bge-base-en-v1.5 0.471 0.997 50.000 48.090 42.310 58.820
BAAI bge-large-en-v1.5 0.404 0.983 49.174 50.380 50.640 51.960
Zephyr Zephyr-7B-beta 0.393 0.981 52.893 46.950 59.620 29.410
Mistral e5-mistral-7b-instruct 0.343 0.971 52.893 48.090 55.770 38.240
Llama 2 Llama-2-7b-hf 0.182 0.964 48.347 44.660 57.690 26.470
Salesforce SFR-Embedding-Mistral 0.343 0.971 50.000 45.420 50.000 40.200
Falcon falcon-7b 0.011 0.964 51.240 48.850 60.900 32.350
Yi Yi-6B Race 0.111 0.964 46.209 64.150 66.170 53.660
Alpaca Alpaca-7B 0.655 0.938 52.357 41.280 41.540 46.340
BAAI bge-base-en-v1.5 0.496 0.992 49.590 44.770 46.250 36.590
BAAI bge-large-en-v1.5 0.404 0.990 50.922 40.890 40.690 51.220
Zephyr Zephyr-7B-beta 0.419 0.992 49.283 42.250 41.330 60.980
Mistral e5-mistral-7b-instruct 0.380 0.999 50.922 52.520 52.680 60.980
Llama 2 Llama-2-7b-hf 0.175 0.990 50.410 45.930 46.680 46.340
Salesforce SFR-Embedding-Mistral 0.381 0.994 51.639 49.030 50.750 39.020
Falcon falcon-7b 0.010 0.985 50.922 46.710 46.900 53.660
Yi Yi-6B Religion 0.147 0.984 52.564 47.620 48.480 33.330
Alpaca Alpaca-7B 0.676 0.823 51.282 80.000 82.830 33.330
BAAI bge-base-en-v1.5 0.497 0.990 46.154 59.050 61.620 16.670
BAAI bge-large-en-v1.5 0.406 0.985 51.282 60.000 61.620 33.330
Zephyr Zephyr-7B-beta 0.465 0.996 51.282 48.570 50.510 16.670
Mistral e5-mistral-7b-instruct 0.436 0.985 52.564 52.380 51.520 66.670
Llama 2 Llama-2-7b-hf 0.202 1.003 44.872 64.760 66.670 33.330
Salesforce SFR-Embedding-Mistral 0.437 0.988 51.282 40.950 39.390 66.670
Falcon falcon-7b 0.009 0.998 48.718 50.480 51.520 33.330
Table 1: Quantitative analysis for DeepSoftDebias using BiasedMAC (BMAC), New SoftMAC (NSMAC), StereotypeScore (SS), Crows Metric Score (CMS), Crows Stereotype Score (CSS), Crows Antistereotype Score (CAS). The best performance is highlighted in bold.
5.2 Stereotype Score
Our research focuses on evaluating and mitigating stereotypical bias in natural language inference tasks using the Stereoset dataset. This dataset comprises pairs of sentences differing only in the substitution of words related to social groups like gender, race, or religion. The objective is to predict their relationship as same, entailment, or contradiction. We introduce a method aimed at reducing bias in word embeddings, with $SS$ values closer to 50 indicating decreased bias. Table 2 presents DeepSoftDebias’s results alongside existing approaches on the Stereoset dataset. Notably, DeepSoftDebias achieves the lowest $SS$ across all social groups, demonstrating its effectiveness in bias reduction. Particularly impressive is DeepSoftDebias’s performance in the gender and race categories, where it significantly outperforms existing methods. For instance, with the SFR-Embedding-Mistral Jiang et al. (2023) model, DeepSoftDebias achieves an $SS$ of 50 for gender and 50.409 for race using the Llama-2-7b model. Additionally, DeepSoftDebias attains a score of 51.282 (or 48.717) for the Zephyr-7b-beta Tunstall et al. (2023) (or Alpaca-7b Taori et al. (2023)) group. We present an illustration of these scores in Fig. 6.
Stereotype Score (SS)
Stereoset Gender Race Religion
FineDeb 53.27 50.82 50.39
CDA 59.61 56.73 58.37
INLP 57.25 57.29 60.31
Self-Debias 59.34 54.30 57.26
Sentence Debias 59.37 57.78 58.73
DeepSoftDebias 50.00 50.41 51.28
Table 2: StereoSet evaluation. Closer to 50 is better for SS. The best performance is highlighted in bold while the next best is underlined).
5.3 Crows-Pairs Dataset
Our study evaluates social bias in natural language generation tasks using the CrowS Pairs dataset, comprising pairs of sentences differing in their degree of bias. By ranking these sentences according to bias level, we quantify the effectiveness of various methods in reducing bias in word embeddings. But as our work is based on word embeddings instead of getting the log-likelihood of the next token from the language model, we compute the average sentence vector for the common parts shared between two sentences. Next, we compare the similarity of this average sentence vector with the uncommon part (i.e., the modified tokens) using word embeddings. By doing so, we capture the semantic differences between stereotypical and non-stereotypical components within the sentence pairs. The rest of the metric remains the same.
Table 3 presents DeepSoftDebias’s results alongside existing approaches on the CrowS Pairs dataset. Notably, DeepSoftDebias achieves scores closest to 50 across all social groups, indicating a significant reduction in social bias. The metric used here is defined in Eq. (2) as follows:
[…]
where stereoscore is the number of stereotypical samples that agree with their label direction and antistereoscore is the number of anti-stereotypical samples that agree with their label direction. Particularly noteworthy is DeepSoftDebias’s superior performance in the gender and religion categories. For instance, with the Yi-6B model, DeepSoftDebias achieves a score of 49.62 for gender and 50.48 for religion with the falcon-7b model. Similarly, using the SFR-Embedding-Mistral model, DeepSoftDebias achieves a score of 49.03 for race biasing the SFR-Embedding-Mistral model. These results underscore the effectiveness of DeepSoftDebias in mitigating social bias in word embeddings. We depict the variation of these scores in Fig. 8.
Crows Pairs Metric Score (CMS)
Crows Pairs Dataset Gender Race Religion
FineDeb 54.58 65.24 44.76
CDA 56.11 56.70 60.00
INLP 51.15 67.96 60.95
Self-Debias 52.29 56.70 56.19
Sentence Debias 52.29 62.72 63.81
DeepSoftDebias 50.38 49.07 50.48
Table 3: Crows Pairs evaluation. Metric score for every demographic. Closer to 50 is better for the metric (best; next best).
6 DownStream Results
Model Variant Bias Direction Biased Text Class. Acc. Debiased Text Class. Acc. Biased NER Macro F1 Avg. Debiased NER Macro F1 Avg.
Mistral e5-mistral-7b-instruct Gender 0.809 0.810 0.385 0.433
Zephyr Zephyr-7B-beta 0.806 0.810 0.368 0.429
Salesforce SFR-Embedding-Mistral 0.807 0.821 0.373 0.428
Yi Yi-6B 0.748 0.782 0.378 0.423
Alpaca Alpaca-7B 0.745 0.729 0.349 0.391
BAAI bge-large-en-v1.5 0.831 0.822 0.456 0.421
BAAI bge-base-en-v1.5 0.831 0.822 0.453 0.392
Mistral e5-mistral-7b-instruct Race 0.811 0.817 0.362 0.435
Zephyr Zephyr-7B-beta 0.806 0.814 0.393 0.443
Salesforce SFR-Embedding-Mistral 0.810 0.818 0.382 0.413
Yi Yi-6B 0.748 0.782 0.378 0.423
Alpaca Alpaca-7B 0.751 0.729 0.358 0.373
BAAI bge-large-en-v1.5 0.832 0.821 0.473 0.424
BAAI bge-base-en-v1.5 0.830 0.813 0.457 0.433
Mistral e5-mistral-7b-instruct Religion 0.808 0.815 0.385 0.437
Zephyr Zephyr-7B-beta 0.805 0.812 0.389 0.444
Salesforce SFR-Embedding-Mistral 0.808 0.817 0.389 0.440
Yi Yi-6B 0.750 0.769 0.384 0.439
Alpaca Alpaca-7B 0.751 0.726 0.364 0.389
BAAI bge-large-en-v1.5 0.830 0.816 0.457 0.419
BAAI bge-base-en-v1.5 0.830 0.817 0.459 0.421
Table 4: Downstream testing results with embeddings debiased using DeepSoftDebias. The first two columns represent results for downstream performance on sentiment analysis. The second two columns represent results for downstream performance on NER. The best performance is highlighted in bold.
Debiasing Direction Biased Baseline Baseline + Adam DeepSoftBias + SGD DeepSoftBias + Adam
Gender 0.390 0.623 0.799 0.893 0.982
Race 0.404 0.656 0.824 0.984 0.987
Religion 0.406 0.623 0.812 0.966 0.983
Table 5: Ablations to characterize various design decisions in the development of DeepSoftDebias. We start with the transformation matrix, then make incremental additions till we reach the proposed architecture of the DeepSoftDebias network.
6.1 Sentiment Classification
In our study, we employ downstream testing to assess the utility of embeddings debiased using DeepSoftDebias across two key natural language processing tasks: text classification and named entity recognition (NER). Utilizing the IMDB Sentiment Classification dataset İlhan Tarımer et al. (2019) for text classification, featuring labeled movie reviews as positive or negative, we compute the average sentence vectors using both original and debiased embeddings. Training XGBoost Chen and Guestrin (2016) classifiers on these vectors, we compare their accuracy on the test set, recognizing accuracy as a straightforward metric for binary classification tasks like sentiment analysis. Notably, our results reveal a performance improvement when debiasing in the gender and religion directions, whereas a slight decrease in performance is observed in the case of race debiasing. We provide a visual representation of these results in Fig. 9.
6.2 Named Entity Recognition (NER)
In our research, we examine the performance of debiased embeddings in the domain of named entity recognition (NER) using the Reuters subset of the CrossNER Liu et al. (2020) dataset. This dataset comprises news domain sentences annotated with four entity types: person, location, organization, and product. Employing a simple BiLSTM model, we input padded arrays of embeddings for each sentence and trained the model on the dataset. We evaluate the models’ performance on the test set using the macro-averaged F1-score, a metric that balances precision and recall, crucial for accurate entity identification and classification. To mitigate potential bias towards more frequent entity types, we adopt macro-averaging, allotting equal importance to each entity type. Remarkably, our findings indicate a slight performance boost when using debiased embeddings in all three directions compared to biased embeddings. We provide a visual representation of these results in Fig. 10.
7 Ablation Experiments
In our study, we conduct ablation experiments to assess the effectiveness of various debiasing techniques in the realm of natural language processing. These techniques encompassed five distinct scenarios: the utilization of debiased embeddings, the application of the original soft debiasing method, the integration of the original debiasing method with the Adam optimizer, the incorporation of new neural networks trained with the debiasing method alongside the SGD optimizer, and finally, our proposed approach, which combines neural networks trained with the debiasing method and the Adam optimizer.
Through rigorous experimentation across three biasing directions, we systematically analyze the performance of each method. Our results reveal a consistent trend of incremental improvements as we transitioned from one method to the next. Notably, DeepSoftDebias, emerged as the standout performer, boasting the highest mean average cosine similarity score across all evaluated scenarios. In addition, our analysis revealed that substituting the transformation matrix with our neural network approach resulted in the most significant enhancement in the efficacy of the debiasing method. This observation underscores the pivotal role played by neural networks in maximizing the effectiveness of the debiasing techniques. We present a visualization of the results of our Ablation experiments in Table 5.
This empirical evidence underscores the robustness and efficacy of our proposed approach in mitigating bias within natural language processing systems. By combining state-of-the-art debiasing techniques with advanced optimization strategies, we have unlocked a powerful methodological framework for enhancing the fairness and accuracy of language models.
8 Discussion
In this section, we summarise the answers to our research questions.
8.1 RQ1
We find that DeepSoftDebias outperforms state-of-the-art methods, and does so without negatively affecting downstream task performance. We make this conclusion after exhaustive testing on several models, and datasets and evaluating several metrics.
8.2 RQ2
We find that size and complexity do affect the ability of debiasing models. Specifically, we make the following observations about DeepSoftDebias:
  • A single layer neural network can effectively de-bias embeddings with $\text{dim}\leq 1024$.
  • A two-layer neural network can effectively debias embeddings with $\text{dim}\leq 2048$.
  • A two-layer neural network with an increased layer size can effectively de-bias embeddings with $\text{dim}\leq 4096$.
  • A three-layer neural network can effectively debias embeddings with $\text{dim}\leq 4450$.
As a step for future work, we are curious to investigate scaling patterns to a further extent.
8.3 RQ3
While debiasing techniques in general can affect the downstream performance of models, we test DeepSoftDebias on multiple challenging downstream tasks and report that our proposed approach, to a large extent, does not negatively influence the performance of different downstream tasks. Remarkably, we see an improvement when using our debiased embeddings for some downstream tasks.
8.4 RQ4
We find that while DeepSoftDebias is effective at reducing bias across gender, race, and religion. We conclude this after testing on multiple embeddings, and multiple datasets and evaluating on multiple performance metrics. As a step for future work, we are curious to investigate whether our proposed approach works towards other forms of bias as well.
9 Conclusion
In this paper, we propose DeepSoftDebias, an approach that leverages neural networks to reduce bias in large language model embeddings. We perform an exhaustive series of tests using multiple performance metrics, state-of-the-art datasets, and downstream tasks to ensure that our debiasing technique is robust, efficient, and accurate. In the future, it would be interesting to see how this method translates to multilingual datasets since bias is language and culture-specific. We hope that this research paves the way for future endeavors that look to make large language models fair, ethical, and bias-free.
10 Limitations
While we do perform exhaustive analysis to test our proposed methodology, our study is monolingual and covers datasets only in English. Consequently, our downstream tasks are also tested only in English. Further, we were unable to test on API-based models at this time.
11 Ethics Statement
We understand that bias can be defined in various ways, and it’s not necessarily ideal for a language model to treat all users exactly the same without considering demographics. There are situations where certain topics require careful handling to avoid perpetuating harmful stereotypes against marginalized communities. Using specific bias metrics might suggest they encompass all negative social impacts across different groups, but we recognize that existing metrics may not capture all nuances in treatment across demographics. Therefore, any benchmark for bias needs to continually evolve to better understand and address these issues as they affect different communities.
The definitions of morality and bias are shaped by cultural perspectives, resulting in diverse interpretations among individuals. Consequently, we do not claim that this work provides an objective or exhaustive measure of any of these concepts.
Appendix
This section provides supplementary material in the form of additional examples, implementation details, etc. to bolster the reader’s understanding of the concepts presented in this work.
Appendix A MAC Scores of DeepSoftDebias
Figures 3, 4, and 5 illustrate how DeepSoftDebias reduces bias in LLM embeddings.
Figure 3: A visual representation of how DeepSoftDebias reduces gender bias in large language model embeddings.
Figure 4: A visual representation of how DeepSoftDebias reduces racial bias in large language model embeddings.
Figure 5: A visual representation of how DeepSoftDebias reduces religion bias in large language model embeddings.
Appendix B Stereoset Scores of DeepSoftDebias
Figures 6 and 8 provide an illustration of word vectors debiased using DeepSoftDebias and their stereoset scores and Crows Metric scores respectively.
Figure 6: A visual representation of word vectors debiased using DeepSoftDebias and their stereotype scores across gender, race and religion respectively.
Figure 7: A visual representation of word vectors debiased using DeepSoftDebias and their Crows Metric score across gender, race and religion respectively.
Figure 8: A visual representation of word vectors debiased using DeepSoftDebias and their Crows Metric scores across gender, race and religion respectively.
Appendix C Downstream Testing Results
Figures 9 and 10 present an illustration of the results of various downstream tasks and their performance evaluation.
Figure 9: An illustration of the results of downstream testing on NER. We compare the performance of biased and debaised embeddings in the directions of gender, race, and religion respectively.
Figure 10: An illustration of results of downstream testing on sentiment analysis. We compare the performance of biased and debaised embeddings in the directions of gender, race, and religion respectively.
Figure 11: An illustration analysis of number of layers in debiasing neural network vs. embedding dimension. We can see the varying performance of the 3 different sizes according to the embedding dimension of the LM it is used with.

### v4

@@
 Embeddings play a pivotal role in the efficacy of large language models.
-They are the bedrock on which these models grasp contextual relationships and foster a more nuanced understanding of language and consequently perform remarkably on a plethora of complex tasks that require a fundamental understanding of human language.
+They are the bedrock on which these models grasp contextual relationships and foster a more nuanced understanding of language and consequently perform complex tasks that require a fundamental understanding of human language.
 Given that these embeddings themselves often reflect or exhibit bias, it stands to reason that these models may also inadvertently learn this bias.
-In this work, we build on the seminal previous work and propose DeepSoftDebias, an algorithm that uses a neural network to perform ‘soft debiasing’.
-We exhaustively evaluate this algorithm across a variety of SOTA datasets, accuracy metrics, and challenging NLP tasks.
+In this work, we build on the aforementioned seminal work of Bolukbasi et al.
+(2016) and Gonen and Goldberg (2019) and propose DeepSoftDebias, an algorithm that uses a neural network to perform ‘soft debiasing’.
+We exhaustively evaluate this algorithm across a variety of state-of-the-art datasets, accuracy metrics, and challenging NLP tasks.
 We find that DeepSoftDebias outperforms the current state-of-the-art methods at reducing bias across gender, race, and religion.
@@
 They act as the basis for these models to understand and subsequently, generate human-like language.
-However, it has been shown that these models rely on word embeddings, which themselves may reflect or exhibit bias Papakyriakopoulos et al.
-(2020); Dev et al.
+However, it has been shown that these word embeddings themselves may reflect or exhibit bias Dev et al.
+(2020); May et al.
+(2019); Caliskan et al.
 (2017).
-Given the exponential increase in the use of large language models on a plethora of downstream tasks, these representations can amplify bias and result in discriminatory actions, especially when it comes to the fields of education, healthcare, and justice.
+Given the exponential increase in the use of LLMs on a plethora of downstream tasks, these representations can amplify bias and result in discriminatory actions, especially when it comes to the fields of education, healthcare, and justice.
 Existing work in this field has looked most commonly into gender bias Kotek et al.
@@
 To answer the above questions, we make the following contributions through this research:
-Our Contributions
-[leftmargin=1mm] We provide, to the best of our knowledge, the first comprehensive study of how various debiasing methods work on SoTA LLM word embeddings We present a novel methodology, DeepSoftDebias, for debiasing LLM word embeddings, which beats SoTA debiasing methods across multiple bias formats including gender, race, and religion.
-We perform an exhaustive quantitative analysis, establishing SoTA baselines and leveraging multiple evaluation metrics to provide a comparison against accessible SoTA baselines
+Our Contributions We provide, to the best of our knowledge, the first comprehensive study of how various debiasing methods work on SoTA LLM word embeddings We present a novel methodology, DeepSoftDebias, for debiasing LLM word embeddings, which beats SoTA debiasing methods across multiple bias formats including gender, race, and religion.
+We perform an exhaustive quantitative analysis, establishing SoTA baselines and leveraging multiple evaluation metrics to provide a comparison against accessible SoTA baselines.
 We illustrate our pipeline in Fig.
@@
 The Debiasing stage leverages the DeepSoftDebias network.
+2 Related Work
+INLP
+Iterative Null-space Projection (INLP) Ravfogel et al.
+(2020) is a post-hoc debiasing method that operates at the representation level.
+The INLP methodology debiases representations by iteratively projecting them into a linear classifier’s null space.
+This technique is particularly effective for handling intersectional groups, which are defined by combinations of sensitive attributes².
+INLP seeks to learn a hidden representation that is independent of the protected attributes.
+This approach is beneficial in scenarios where an attempt to make a model fairer towards some group results in increased unfairness towards another group.
+Therefore, INLP emerges as a robust and effective strategy for mitigating bias in language models, promoting fairness across multiple protected attributes.
+Self-Debias
+Self-Debiasing Schick et al.
+(2021) is a novel approach to mitigating bias in language models.
+The methodology, first coined by Schick et al.
+(2021), is based on the concept of self-diagnosis.
+In this approach, pretrained language models recognize their undesirable biases and the toxicity of the content they produce.
+Based on this self-diagnosis, a decoding algorithm is proposed that reduces the probability of a language model producing problematic text.
+This approach, referred to as self-debiasing, does not rely on manually curated word lists, nor does it require any training data or changes to the model’s parameters.
+While it does not completely eliminate the issue of language models generating biased text, it is an important step in this direction.
+The self-debiasing approach demonstrates the potential of language models to self-regulate and reduce their inherent biases.
+Sentence Debias
+SentenceDebias Liang et al.
+(2020) is a debiasing methodology that operates at the sentence level.
+It is a projection-based method that identifies a linear subspace associated with a specific bias.
+The sentence representations are projected onto this bias subspace, and the projection is subtracted from the original representations.
+This process effectively debiases the sentence representations.
+SDB is particularly useful for mitigating biases related to gender, race, and religion.
+It offers a comprehensive comparison between models that adjust weights for debiasing and those employing test-time surgical interventions.
+The SDB method signifies a significant advancement in debiasing strategies, promoting a more equitable representation in language models.
+Counterfactual Data Augumentation
+Counterfactual Data Augmentation (CDA) Yadav et al.
+(2023) is a data-based debiasing strategy often used to mitigate gender bias.
+The CDA methodology involves re-balancing a corpus by swapping bias attribute words (e.g., he/she) in a dataset.
+This technique is part of a broader set of debiasing techniques that also includes Dropout, Self-Debias, SentenceDebias, and Iterative Nullspace Projection.
+CDA has been applied to various language models, including BERT, with the goal of diminishing stereotypical biases while maintaining the model’s performance on downstream tasks.
+However, it’s important to note that while CDA has the potential to improve the fairness of NLP models, it may not be effective in eliminating all biases and may even introduce new biases or errors in the model³.
+FineDeb
+FineDeb Saravanan et al.
+(2023) is a two-phase debiasing framework for language models.
+In the first phase, FineDeb debiases the model by modifying the embeddings learned by the language model.
+This process involves contextual debiasing of these embeddings.
+In the second phase, the debiased model is fine-tuned on the language modeling objective.
+This methodology is effective for demographics with multiple classes.
+The FineDeb approach demonstrates its effectiveness through extensive experiments and comparisons with state-of-the-art techniques.
+It offers stronger debiasing in comparison to other methods, which often result in models as biased as the original language model.
+Thus, FineDeb emerges as a robust and effective framework for mitigating bias in language models.
 3 Data
 This study leverages several datasets to examine and address biases in word embeddings and language models, focusing on the representation and perpetuation of stereotypes within these systems.
-2.1 L2-Reddit Corpus
+L2-Reddit Corpus
 We utilize the L2-Reddit [^1] https://github.com/ellarabi/reddit-l2 Rabinovich et al.
-(2018) corpus, a collection of Reddit posts and comments by both native and non-native English speakers, featuring approximately 56 million sentences from the United States.
-This dataset serves as the foundation for training the word embeddings, aiming to capture the nuanced and inherently biased linguistic patterns present in social media discourse.
-The corpus provides a rich linguistic resource for analyzing and understanding the initial biased word embeddings.
+(2018) corpus, a collection of Reddit posts and comments by both native and non-native English speakers, featuring approximately 56 million sentences.
+This dataset serves as the foundation for training word embeddings, aiming to capture the nuanced and inherently biased linguistic patterns present in social media discourse.
 In our study, we employ the Reddit L2 corpus as the source for our initial Word2Vec Mikolov et al.
 (2013) word embeddings.
-Subsequently, we leverage the vocabulary derived from these word vectors to obtain the word embeddings from the Language Model Models (LLMs).
-2.2 StereoSet
+Subsequently, we leverage the vocabulary derived from these word vectors to obtain the word embeddings from the LLMs.
+We utilize Word2Vec on the Reddit-L2 corpus to obtain the vocabulary.
+This vocabulary comprises the words for which we aim to extract embeddings from the LLMs.
+The primary objective of this approach is to ensure a consistent set of words across all our LLMs.
+This consistency allows each of our LLMs to be tested on the same set of words.
+StereoSet
 StereoSet Nadeem et al.
@@
 (2020).
-The dataset’s creation involved a rigorous crowdsourcing process, ensuring a wide coverage of stereotypes and anti-stereotypes to challenge and evaluate the fairness of NLP systems.
-In our words we use the stereoset dataset to benchmark our debiasing method against all the other debiasing methods.
-To accomplish this, we compute the average sentence vector for all sentences and their respective contexts.
-We then compare these vectors against each other to derive the stereotype score.
-2.3 CrowS-Pairs
+In our work, we use the Stereoset dataset to benchmark our debiasing method.
+CrowS-Pairs
 CrowS-Pairs Nangia et al.
@@
 It contrasts sentences related to historically disadvantaged and advantaged groups in the U.S., with annotations from crowd workers highlighting the degree of stereotyping.
-Despite its noise and reliability issues, CrowS-Pairs remains a valuable resource for evaluating the bias in major language models, providing a nuanced metric that quantifies preference for stereotypical versus non-stereotypical expressions.
 In our study, we obtain debiased word embeddings for sentences by computing the average sentence vector for both less and more stereotypical or anti-stereotypical directions.
@@
 4 Methodology
-In this research, we delve into the domain of debiasing word embeddings, presenting both an established and a newly proposed methodology for mitigating biases in word vector representations.
+In this section, we delve into the domain of debiasing word embeddings, presenting both an established and a newly proposed methodology for mitigating biases in word vector representations.
 These biases span across gender, racial, and religious lines and are encoded inadvertently within language models.
 4.1 Bias Identification and Data Structure
-To quantitatively assess bias in word embeddings, we measure the projection of word vectors onto a gender-specific axis, defined by the vectorial difference between the terms ‘he’ and ‘she.’ The magnitude of this projection serves as an indicator of bias.
+To quantitatively assess bias in word embeddings, we measure the projection of word vectors onto a gender-specific axis, defined by the vector difference between the terms ‘he’ and ‘she.’ The magnitude of this projection serves as an indicator of bias.
 We use a structured vocabulary with its associated vector representations from the Word2Vec model to facilitate the identification of biases.
@@
 These resources are crucial for the systematic identification and rectification of biases in word embeddings.
+The words used for the BiasSpace are present in AppendixA.
 4.2 Soft Debiasing: The Baseline Approach
-The initial method as seen in the paper Manzini et al.
+The initial method as seen in Manzini et al.
 (2019) leverages a method called soft debiasing.
 We recap its algorithm in Algorithm 1.
-Soft debiasing involves learning a projection of the embedding matrix that preserves the inner product between biased and debiased embeddings while minimizing the projection onto the bias subspace of embeddings mentioned in 3.1 that should be neutral.
-Given embeddings $W$ and $N$ which are embeddings for the whole vocabulary and the subset of bias-neutral words respectively, and the bias subspace B obtained in Section 2.1, soft debiasing seeks a linear transformation A that minimizes the following objective defined in Eq.
+Soft debiasing involves learning a projection of the embedding matrix that preserves the inner product between biased and debiased embeddings while minimizing the projection onto the bias subspace of embeddings mentioned in 4.1.
+Given embeddings $W$ and $N$ which are embeddings for the whole vocabulary and the subset of bias-neutral words respectively, and the bias subspace B obtained in Section 3, soft debiasing seeks a linear transformation A that minimizes the following objective defined in Eq.
 (1) as follows:
@@
 Minimizing the first term preserves the inner product after the linear transformation A, and minimizing the second term minimizes the projection onto the bias subspace B of embeddings.
-R is a tunable parameter that balances the two objectives.
+$\lambda$ is a tunable parameter that balances the two objectives.
+W here refers to the matrix of word embeddings and N refers to the matrix of the embeddings of the neutral space i.e.
+words that aren’t influenced by any bias.
 Input: Biased word embeddings ($\text{emb}_{\text{biased}}$), Bias Subspace (BiasSpace), Neutral word embeddings ($\text{emb}_{\text{neutral}}$)
@@
 We demonstrate the effectiveness of DeepSoftDebias on various datasets and tasks, and show that it outperforms the state-of-the-art methods in terms of accuracy and efficiency.
-Fig.
+The reason for the need of a fixed Biasspace is that we adopt the methodology proposed by Bolukbasi et al.
+for the derivation of the bias subspace.
+The process of creating the BiasSpace commences with the identification of word vectors representing opposing concepts, such as ‘he’ versus ‘she’, or ‘man’ versus ‘woman’.
+For each pair, we compute the mean vector, which encapsulates the shared semantic space.
+Subsequently, we subtract this mean vector from the original word vectors, yielding vectors that exclusively represent the bias components.
+These bias vectors are then concatenated to form a matrix, referred to as the bias subspace.
+This bias subspace plays a pivotal role in the training of our neural network.
+Specifically, we ensure that the output of the word embeddings, upon being processed through the neural network, is orthogonal to the bias subspace Fig.
 2 presents a visualization of our approach to downstream testing.
@@
 Algorithm 2 Neural Network Approach
-Figure 2: A step-by-step visualization of our downstream testing process to effectively evaluate DeepSoftDebias
+Figure 2: A step-by-step visualization of our downstream testing process to effectively evaluate DeepSoftDebias.
 5 Effects of LLM Size and Dependency of Network Size
@@
 (2023) and Yi-6b 01.ai (2024) need a more complex debiasing neural network.
-For embedding dimensions around 2000, a two-layer neural network is sufficient, while for larger embedding dimensions, a three-layer neural network is required to achieve good debiasing results.
+For embeddings with embedding length of around 2000, a two-layer neural network is sufficient, while for larger embedding dimensions, a three-layer neural network is required to achieve good debiasing results.
+In addressing the second research question, we delve into the intricacies of neural network complexity necessary for debiasing embeddings of varying sizes.
+While our discussion highlights the effectiveness of larger neural networks in mitigating bias within Language Model (LM) embeddings with substantial dimensions, it is imperative to substantiate this observation.
+We would like to point out that we draw inspiration from the conceptual framework of DeepSoftDebias.
+Building upon the foundational work by Bolukbasi et al., which employed a transformation matrix for word embedding debiasing, our approach replaces this matrix with a neural network.
+This neural network can be conceptualized as a series of interconnected matrices.
+Specifically, when de-biasing larger LMs with embedding dimensions exceeding 4096, we augment the neural network by increasing the number of layers and adjusting layer sizes.
+This augmentation enables us to model the intricate dependencies inherent in debiasing processes for larger embedding dimensions.
+Consequently, deeper neural networks emerge as more efficacious tools for addressing bias in such expansive models.
 Additionally, the debiasing neural network and the optimization algorithm need to be hyperparameter-tuned, such as adjusting the learning rate, to get optimal results.
@@
 6 Results
-The results obtained are described in the following subsections.
-5.1 Mean Average Cosine Similarity (MAC)
-MAC Manzini et al.
-(2019) is a metric used to quantify semantic associations between word classes and attributes.
-MAC takes word embeddings, targets (representing classes), and attributes as inputs.
-By computing the mean cosine distance between target words and attribute sets, MAC offers a concise measure of semantic proximity.
-This metric provides valuable insights into the contextual semantics encoded within word embeddings.
-Table 1 shows that the word embeddings debiased in the direction of race and gender have comparable increases in their average MAC of  0.64, whereas word embeddings debiased in the direction of religion have an increase in MAC of around 0.61.
-We see that our debiasing procedure categorically moves MAC scores closer to 1.0.
-This indicates an increase in cosine distance.
-Further, the associated P-values indicate these changes are statistically significant.
-This demonstrates that our approach for multiclass debiasing decreases bias in the word embeddings.
-We provide visual representations of the efficiency of DeepSoftDebias at removing gender bias, racial bias, and religion bias in Appendix A.
+In this section, we provide an extensive analysis of our proposed methodology, complete with a comprehensive evaluation against multiple metrics, tasks, and datasets.
+We provide the results of additional downstream testing and ablation experiments in Appendix D and Appendix F, respectively.
+We also provide our hypothesis of why there is a variation in bias across LLMs in Appendix E.
 Model Variant Topic BMAC NSMAC SS CMS CSS CAS
@@
 Falcon falcon-7b 0.011 0.964 51.240 48.850 60.900 32.350
+Gemma Gemma-2b 0.058 0.971 47.107 48.470 57.69 36.27
+Gemma Gemma-7b 0.553 0.976 49.173 51.53 63.46 35.29
+GritLM GritLM-7B 0.379 0.999 51.239 48.470 57.05 37.25
+mxbai mxbai-embed-large-v1 0.467 0.994 51.652 55.34 60.9 49.02
 Yi Yi-6B Race 0.111 0.964 46.209 64.150 66.170 53.660
@@
 The best performance is highlighted in bold.
+6.1 Mean Average Cosine Similarity
+Mean Average Cosine Similarity (MAC) Manzini et al.
+(2019) is a metric used to quantify semantic associations between word classes and attributes.
+MAC takes word embeddings, targets (representing classes), and attributes as inputs.
+By computing the mean cosine distance between target words and attribute sets, MAC offers a concise measure of semantic proximity.
+This metric provides valuable insights into the contextual semantics encoded within word embeddings.
+Table 1 shows that the word embeddings debiased in the direction of race and gender have comparable increases in their average MAC of 0.64, whereas word embeddings debiased in the direction of religion have an increase in MAC of 0.61.
+We see that our debiasing procedure categorically moves MAC scores closer to 1.0.
+This indicates an increase in cosine distance.
+Further, the associated P-values indicate these changes are statistically significant.
+This demonstrates that our approach for multiclass debiasing decreases bias in the word embeddings.
+We provide visual representations of the efficiency of DeepSoftDebias at removing gender bias, racial bias, and religion bias in Appendix B.
+In our research, we have chosen to utilize Mean Average Cosine Similarity (MAC) as our primary metric for assessing bias in word embeddings.
+This decision is informed by the work of Manzini et al.
+(2019), who posit that MAC can be viewed as an extension of the Word Embedding Association Test (WEAT), specifically adapted for a multiclass setting.
+The MAC and WEAT serve distinct, yet complementary purposes.
+While WEAT is designed to focus on specific associations between word vectors and predefined concepts (such as gender or race), MAC provides a broader perspective by measuring overall similarity patterns across different groups.
+This makes MAC less sensitive to specific word choices, thereby revealing biases that might be overlooked by WEAT.
+In essence, both metrics contribute to a comprehensive understanding of bias in word embeddings.
+However, the use of MAC is particularly beneficial in our research as it complements the findings of WEAT, providing a more holistic view of bias in the data.
+This approach allows us to capture a wider range of biases, thereby enhancing the robustness of our analysis.
 6.2 Stereotype Score
-Our research focuses on evaluating and mitigating stereotypical bias in natural language inference tasks using the Stereoset dataset.
+Our research focuses on evaluating and mitigating stereotypical bias in NLI tasks using the Stereoset dataset.
 This dataset comprises pairs of sentences differing only in the substitution of words related to social groups like gender, race, or religion.
 The objective is to predict their relationship as same, entailment, or contradiction.
-We introduce a method aimed at reducing bias in word embeddings, with $SS$ values closer to 50 indicating decreased bias.
+We introduce a method aimed at reducing bias in word embeddings, with Stereotype Score $SS$ values closer to 50 indicating decreased bias.
 Table 2 presents DeepSoftDebias’s results alongside existing approaches on the Stereoset dataset.
@@
 (2023) model, DeepSoftDebias achieves an $SS$ of 50 for gender and 50.409 for race using the Llama-2-7b model.
-Additionally, DeepSoftDebias attains a score of 51.282 (or 48.717) for the Zephyr-7b-beta Tunstall et al.
-(2023) (or Alpaca-7b Taori et al.
-(2023)) group.
-We present an illustration of these scores in Fig.
+Additionally, DeepSoftDebias attains a score of 51.282 for the Zephyr-7b-beta Tunstall et al.
+(2023) or 48.717 for Alpaca-7b Taori et al.
+(2023).
+We present these score in 2 an illustration of these scores in Fig.
 6.
@@
 […]
-where stereoscore is the number of stereotypical samples that agree with their label direction and antistereoscore is the number of anti-stereotypical samples that agree with their label direction.
+where Crows Pair Stereotype Score (CSS) is the number of stereotypical samples that agree with their label direction and Crows Pairs Anti-stereotype Score (CAS) is the number of anti-stereotypical samples that agree with their label direction.
+Label direction refers to the label given the pair of sentences whether they are stereotypical or anti-stereotypical.
+In our evaluation we get the average sentence vector of the context and the more and less (anti-)stereotypical sentence.
+We then see whether the context vector is closer to the more (anti-)stereotypical sentence or the less (anti-)stereotypical sentence.
+If it is closer to the more (anti–)stereotypical sentence,, then we state that it agrees with the (anti–)stereotype, i.e., the label direction.
 Particularly noteworthy is DeepSoftDebias’s superior performance in the gender and religion categories.
@@
 These results underscore the effectiveness of DeepSoftDebias in mitigating social bias in word embeddings.
-We depict the variation of these scores in Fig.
+We present these score in 3 and depict the variation of these scores in Fig.
 8.
+We also report the CSS and CAS score which refer to the CrossNER Stereotype score, i.e., the number of times the model agrees with the more stereotypes statement when the label direction is stereotype, and the CrossNER Anti-stereotype score, which refers to the number of times the model agrees with the more anti-stereotyped statement when the label direction was anti-sterotype.
 Crows Pairs Metric Score (CMS)
@@
 Table 3: Crows Pairs evaluation. Metric score for every demographic. Closer to 50 is better for the metric (best; next best).
-6 DownStream Results
+7 Discussion
+In this section, we summarise the answers to our research questions.
+RQ1
+We find that DeepSoftDebias outperforms state-of-the-art methods, and does so without negatively affecting downstream task performance.
+We make this conclusion after exhaustive testing on several models, and datasets and evaluating several metrics.
+RQ2
+We find that size and complexity do affect the ability of debiasing models. Specifically, we make the following observations about DeepSoftDebias:
+  • A single layer neural network can effectively de-bias embeddings with $\text{dim}\leq 1024$.
+  • A two-layer neural network can effectively debias embeddings with $\text{dim}\leq 2048$.
+  • A two-layer neural network with an increased layer size can effectively de-bias embeddings with $\text{dim}\leq 4096$.
+  • A three-layer neural network can effectively debias embeddings with $\text{dim}\leq 4450$.
+As a step for future work, we are curious to investigate scaling patterns to a further extent.A visualization of this is provided in Fig 11
+RQ3
+While debiasing techniques in general can affect the downstream performance of models, we test DeepSoftDebias on multiple challenging downstream tasks and report that our proposed approach, to a large extent, does not negatively influence the performance of different downstream tasks.
+Remarkably, we see an improvement when using our debiased embeddings for some downstream tasks.
+RQ4
+We find that while DeepSoftDebias is effective at reducing bias across gender, race, and religion.
+We conclude this after testing on multiple embeddings, and multiple datasets and evaluating on multiple performance metrics.
+As a step for future work, we are curious to investigate whether our proposed approach works towards other forms of bias as well.
+8 Conclusion
+In this paper, we propose DeepSoftDebias, an approach that leverages neural networks to reduce bias in large language model embeddings.
+We perform an exhaustive series of tests using multiple performance metrics, state-of-the-art datasets, and downstream tasks to ensure that our debiasing technique is robust, efficient, and accurate.
+In the future, it would be interesting to see how this method translates to multilingual datasets since bias is language and culture-specific.
+We hope that this research paves the way for future endeavors that look to make LLMs fair, ethical, and bias-free.
+9 Limitations
+While we do perform exhaustive analysis to test our proposed methodology, our study is monolingual and covers datasets only in English.
+Consequently, our downstream tasks are also tested only in English.
+Further, we were unable to conduct test on API-based models at this time.
+Our testing was also constrained by the limitations of GPU VRAM, which prevented us from extending our testing to larger models such as Llama-65B.
+These models could not be accommodated within the GPU VRAM, even after applying quantization to 8 bits.
+Consequently, the largest model that we were able to test was the Gemma-7B model.
+10 Ethics Statement
+We understand that bias can be defined in various ways, and it’s not necessarily ideal for a language model to treat all users exactly the same without considering demographics.
+There are situations where certain topics require careful handling to avoid perpetuating harmful stereotypes against marginalized communities.
+Using specific bias metrics might suggest they encompass all negative social impacts across different groups, but we recognize that existing metrics may not capture all nuances in treatment across demographics.
+Therefore, any benchmark for bias needs to continually evolve to better understand and address these issues as they affect different communities.
+The definitions of morality and bias are shaped by cultural perspectives, resulting in diverse interpretations among individuals.
+Consequently, we do not claim that this work provides an objective or exhaustive measure of any of these concepts.
+Appendix
+This section provides supplementary material in the form of additional examples, implementation details, etc. to bolster the reader’s understanding of the concepts presented in this work.
+Appendix A Table of words and bias they represent
+Bias Direction Biased Words
+Gender Male "manager", "executive", "doctor", "lawyer", "programmer", "scientist", "soldier", "supervisor", "rancher", "janitor", "firefighter", "officer"
+Female "secretary", "nurse", "clerk", "artist", "homemaker", "dancer", "singer", "librarian", "maid", "hairdresser", "stylist", "receptionist", "counselor"
+Race Black "slave", "musician", "runner", "criminal", "homeless"
+Asian "manager", "executive", "redneck", "hillbilly", "leader", "farmer"
+Caucasian "doctor", "engineer", "laborer", "teacher"
+Religion Jew "greedy", "cheap", "hairy", "liberal"
+Christian "judgemental", "conservative", "familial"
+Muslim "violent", "terrorist", "dirty", "uneducated"
+Table 4: List of Words related to sub-categories in the bias directions explored
+Appendix B MAC Scores of DeepSoftDebias
+Figures 3, 4, and 5 illustrate how DeepSoftDebias reduces bias in LLM embeddings.
+Figure 3: A visual representation of how DeepSoftDebias reduces gender bias in large language model embeddings.
+Figure 4: A visual representation of how DeepSoftDebias reduces racial bias in large language model embeddings.
+Figure 5: A visual representation of how DeepSoftDebias reduces religion bias in large language model embeddings.
+Appendix C Stereoset Scores of DeepSoftDebias
+Figures 6 and 8 provide an illustration of word vectors debiased using DeepSoftDebias and their stereoset scores and Crows Metric scores respectively.
+Figure 6: A visual representation of word vectors debiased using DeepSoftDebias and their stereotype scores across gender, race and religion respectively.
+Figure 7: A visual representation of word vectors debiased using DeepSoftDebias and their Crows Metric score across gender, race and religion respectively.
+Figure 8: A visual representation of word vectors debiased using DeepSoftDebias and their Crows Metric scores across gender, race and religion respectively.
+Appendix D Downstream Testing Results
+Model Variant Baseline Debiased Text Class. Acc. DSB Debiased Text Class. Acc. Baseline Debiased NER Macro F1 Avg. DSB Debiased NER Macro F1 Avg.
+Gemma gemma-2b 0.7655 0.7964 0.469 0.484
+BAAI bge-base-en-v1.5 0.8296 0.822 0.458 0.421
+Mistral SFR-Embedding-Mistral 0.8297 0.821 0.404 0.428
+Gemma gemma-7b 0.516 0.8032 0.198 0.475
+Zephyr zephyr-7b-beta 0.7997 0.81 0.403 0.429
+mxbai mxbai-embed-large-v1 0.8366 0.7903 0.461 0.455
+Table 5: Downstream testing results comparison with embeddings debiased using DeepSoftDebias and the baseline SoftDebais Method.
+The first two columns represent results for downstream performance on sentiment analysis.
+The second two columns represent results for downstream performance on NER.
+In our research, we primarily focus on the debiasing of word embeddings derived from Language Learning Models (LLMs).
+We aim to investigate the impact of this debiasing on the performance of these embeddings when subjected to identical training and testing methodologies.
+Our objective is to quantitatively measure any performance fluctuations (increase or decrease) on the downstream tasks that we test.
+For this purpose, we trained simple models on top of these word embeddings.
+For instance, we used an XGBoost model without any hyperparameter tuning for the classification task, and a straightforward bidirectional LSTM for the Named Entity Recognition (NER) task.
+It is important to note that our goal in presenting our results on these two tasks is not to establish a benchmark for debiased embeddings.
+Instead, we aim to demonstrate the effect of debiasing on the performance of word embeddings in downstream tasks, as seen in the seminal work of Gonen and Goldberg (2019).
+This approach allows us to provide a more comprehensive understanding of the implications and potential benefits of debiasing word embeddings.
+D.1 Sentiment Classification
+In our study, we employ downstream testing to assess the utility of embeddings debiased using DeepSoftDebias across two key natural language processing tasks: text classification and named entity recognition (NER).
+Utilizing the IMDB Sentiment Classification dataset İlhan Tarımer et al.
+(2019) and Stanford TreeBank Dataset for text classification, featuring labeled movie reviews as positive or negative, we compute the average sentence vectors using both original and debiased embeddings.
+Training XGBoost Chen and Guestrin (2016) classifiers on these vectors, we compare their accuracy on the test set, recognizing accuracy as a straightforward metric for binary classification tasks like sentiment analysis.
+Notably, our results reveal a performance improvement when debiasing in the gender and religion directions, whereas a slight decrease in performance is observed in the case of race debiasing.
+We provide these results in Table:6 for IMDB Sentiment classification and Table:7 for Stanford Sentiment Treebank.
+A visual representation of these results in Fig.
+9.
+D.2 Named Entity Recognition (NER)
+In our research, we examine the performance of debiased embeddings in the domain of named entity recognition (NER) using the Reuters subset of the CrossNER Liu et al.
+(2020) dataset.
+This dataset comprises news domain sentences annotated with four entity types: person, location, organization, and product.
+Employing a simple BiLSTM model, we input padded arrays of embeddings for each sentence and trained the model on the dataset.
+We evaluate the models’ performance on the test set using the macro-averaged F1-score, a metric that balances precision and recall, crucial for accurate entity identification and classification.
+To mitigate potential bias towards more frequent entity types, we adopt macro-averaging, allotting equal importance to each entity type.
+Remarkably, our findings indicate a slight performance boost when using debiased embeddings in all three directions compared to biased embeddings.
+We provide these results in Table:6and a visual representation of these results in Fig.
+10.
+D.3 Semantic Textual Similarity
+In our research, we evaluate the performance of debiased embeddings for the Semantic Textual Similarity (STS) task using the STS-B dataset.
+This dataset, a component of the General Language Understanding Evaluation (GLUE) benchmark, is a valuable resource for the STS task.
+The task aims to quantify the semantic similarity between two sentences, assigning a score from 1 to 5 based on their degree of semantic equivalence.
+The STS-B dataset, comprising examples from diverse sources, includes human annotations for sentence pair similarity, contributing significantly to the broader field of natural language understanding by facilitating the measurement of meaning equivalence across sentences.
+To utilize the embeddings for the task, we train a dual-head neural network.
+We perform cosine similarity after passing the average sentence vector of the two sentences through the network, followed by a Fully Connected layer to obtain the actual score.
+The performance of our approach is evaluated using Pearson’s correlation and Spearman’s correlation as metrics.
+This methodology allows us to develop and evaluate models’ ability to understand nuanced semantic relationships in text effectively.
+We provide our results in this task in Table:7
+Figures 9 and 10 present an illustration of the results of various downstream tasks and their performance evaluation.
+Figure 9: An illustration of the results of downstream testing on NER. We compare the performance of biased and debaised embeddings in the directions of gender, race, and religion respectively.
+Figure 10: An illustration of results of downstream testing on sentiment analysis.
+We compare the performance of biased and debaised embeddings in the directions of gender, race, and religion respectively.
+Figure 11: An illustration analysis of number of layers in debiasing neural network vs.
+embedding dimension.
+We can see the varying performance of the 3 different sizes according to the embedding dimension of the LM it is used with.
 Model Variant Bias Direction Biased Text Class. Acc. Debiased Text Class. Acc. Biased NER Macro F1 Avg. Debiased NER Macro F1 Avg.
@@
 The best performance is highlighted in bold.
+Model Name Topic STS-B $\uparrow$ Baseline Debiased PCC STS-B $\uparrow$ DeepSoftDebias Debiased PCC SST Biased Acc. SST Baseline Debiased Acc. SST DeepSoftDebias Debiased Acc.
+BAAI/bge-base-en-v1.5 Gender 0.088 0.001 0.730 0.725 0.693
+BAAI/bge-large-en-v1.5 0.159 0.105 0.727 0.710 0.705
+google/gemma-2b -0.060 0.154 0.686 0.677 0.678
+google/gemma-7b -0.059 0.017 0.675 0.544 0.691
+GritLM/GritLM-7B -0.125 0.044 0.711 0.702 0.697
+HuggingFaceH4/zephyr-7b-beta -0.129 0.097 0.706 0.687 0.699
+intfloat/multilingual-e5-large-instruct -0.037 0.096 0.729 0.720 0.724
+meta-llama/Llama-2-7b-hf 0.009 -0.032 0.701 0.692 0.686
+openai-community/gpt2-large 0.042 -0.038 0.664 0.665 0.669
+openai-community/gpt2-xl 0.041 0.071 0.666 0.667 0.669
+tiiuae/falcon-7b -0.116 0.066 0.686 0.672 0.694
+BAAI/bge-base-en-v1.5 Race 0.094 0.092 0.730 0.709 0.683
+BAAI/bge-large-en-v1.5 0.104 0.099 0.727 0.727 0.695
+google/gemma-2b -0.041 0.164 0.686 0.665 0.686
+google/gemma-7b -0.055 0.133 0.675 0.549 0.678
+GritLM/GritLM-7B -0.133 -0.057 0.711 0.714 0.690
+HuggingFaceH4/zephyr-7b-beta -0.127 0.062 0.706 0.687 0.697
+intfloat/multilingual-e5-large-instruct 0.053 0.120 0.729 0.730 0.730
+meta-llama/Llama-2-7b-hf -0.058 0.113 0.701 0.699 0.705
+openai-community/gpt2-large -0.019 0.024 0.664 0.670 0.680
+openai-community/gpt2-xl 0.149 0.180 0.666 0.665 0.692
+tiiuae/falcon-7b -0.192 -0.027 0.686 0.664 0.693
+BAAI/bge-base-en-v1.5 Religion 0.054 0.078 0.730 0.716 0.694
+BAAI/bge-large-en-v1.5 0.153 0.175 0.727 0.718 0.697
+google/gemma-2b 0.118 0.278 0.686 0.679 0.682
+google/gemma-7b 0.127 0.194 0.675 0.548 0.685
+GritLM/GritLM-7B -0.002 0.077 0.711 0.702 0.703
+HuggingFaceH4/zephyr-7b-beta -0.130 0.118 0.706 0.693 0.686
+intfloat/multilingual-e5-large-instruct 0.201 0.194 0.729 0.728 0.735
+meta-llama/Llama-2-7b-hf -0.103 0.032 0.701 0.679 0.710
+openai-community/gpt2-xl 0.247 0.251 0.666 0.671 0.679
+tiiuae/falcon-7b 0.126 0.265 0.686 0.671 0.703
+Table 7: Downstream testing results on Stanford Sentiment Treebank and STS-B Semantic Similarity Dataset.
+PCC here refers to the Pearson’s Coefficient and we report the gain in positive PCC from the Biased embeddings to the debiased embeddings.
+SST is Stanford Sentiment Treebank and STS-B is the Semantic Textual Similarity Benchmark
+Appendix E Variation of Bias in the Different LLMs
+The presence of biases in has drawn significant attention from researchers and practitioners.
+These biases can inadvertently emerge during the training process due to the characteristics of the initial training data.
+In this study, we explore the factors contributing to bias variation among LLMs, focusing on three prominent models: Llama, Mistral, and Gemma.
+Our analysis reveals that biases, including those related to gender, race, and culture, are often inherited from the training data.
+For instance, historical texts may perpetuate gender stereotypes or racial prejudices present in their source material.
+Llama and Mistral, trained on diverse corpora containing web documents, source code, and mathematical text, exhibit varying degrees of bias.
+Gemma, released by Google, further demonstrates the impact of training data size, with both 2B and 7B variants drawing from an extensive pool of up to 6 trillion tokens.
+Appendix F Ablation Experiments
+In our study, we conduct ablation experiments to assess the effectiveness of various debiasing techniques in the realm of natural language processing.
+These techniques encompassed five distinct scenarios: the utilization of debiased embeddings, the application of the original soft debiasing method, the original debiasing method with the Adam optimizer, DeepSoftDebias with the SGD optimizer, and finally, DeepSoftDebias with the Adam optimizer.
+These experiments were gauged based on MAC as the evaluation metric.
+Through rigorous experimentation across three biasing directions, we systematically analyze the performance of each method.
+Our results reveal a consistent trend of incremental improvements as we transitioned from one method to the next.
+Notably, DeepSoftDebias, emerged as the standout performer, boasting the highest mean average cosine similarity score across all evaluated scenarios.
+In addition, our analysis revealed that substituting the transformation matrix with our neural network approach resulted in the most significant enhancement in the efficacy of the debiasing method.
+This observation underscores the pivotal role played by neural networks in maximizing the effectiveness of the debiasing techniques.
+Table 8 presents a visualization of the results of our ablation experiments.
+This empirical evidence underscores the robustness and efficacy of our proposed approach in mitigating bias within natural language processing systems.
+By combining state-of-the-art debiasing techniques with advanced optimization strategies, we have unlocked a powerful methodological framework for enhancing the fairness and accuracy of language models.
 Debiasing Direction Biased Baseline Baseline + Adam DeepSoftBias + SGD DeepSoftBias + Adam
@@
 We start with the transformation matrix, then make incremental additions till we reach the proposed architecture of the DeepSoftDebias network.
-6.1 Sentiment Classification
-In our study, we employ downstream testing to assess the utility of embeddings debiased using DeepSoftDebias across two key natural language processing tasks: text classification and named entity recognition (NER).
-Utilizing the IMDB Sentiment Classification dataset İlhan Tarımer et al.
-(2019) for text classification, featuring labeled movie reviews as positive or negative, we compute the average sentence vectors using both original and debiased embeddings.
-Training XGBoost Chen and Guestrin (2016) classifiers on these vectors, we compare their accuracy on the test set, recognizing accuracy as a straightforward metric for binary classification tasks like sentiment analysis.
-Notably, our results reveal a performance improvement when debiasing in the gender and religion directions, whereas a slight decrease in performance is observed in the case of race debiasing.
-We provide a visual representation of these results in Fig.
-9.
-6.2 Named Entity Recognition (NER)
-In our research, we examine the performance of debiased embeddings in the domain of named entity recognition (NER) using the Reuters subset of the CrossNER Liu et al.
-(2020) dataset.
-This dataset comprises news domain sentences annotated with four entity types: person, location, organization, and product.
-Employing a simple BiLSTM model, we input padded arrays of embeddings for each sentence and trained the model on the dataset.
-We evaluate the models’ performance on the test set using the macro-averaged F1-score, a metric that balances precision and recall, crucial for accurate entity identification and classification.
-To mitigate potential bias towards more frequent entity types, we adopt macro-averaging, allotting equal importance to each entity type.
-Remarkably, our findings indicate a slight performance boost when using debiased embeddings in all three directions compared to biased embeddings.
-We provide a visual representation of these results in Fig.
-10.
-7 Ablation Experiments
-In our study, we conduct ablation experiments to assess the effectiveness of various debiasing techniques in the realm of natural language processing.
-These techniques encompassed five distinct scenarios: the utilization of debiased embeddings, the application of the original soft debiasing method, the integration of the original debiasing method with the Adam optimizer, the incorporation of new neural networks trained with the debiasing method alongside the SGD optimizer, and finally, our proposed approach, which combines neural networks trained with the debiasing method and the Adam optimizer.
-Through rigorous experimentation across three biasing directions, we systematically analyze the performance of each method.
-Our results reveal a consistent trend of incremental improvements as we transitioned from one method to the next.
-Notably, DeepSoftDebias, emerged as the standout performer, boasting the highest mean average cosine similarity score across all evaluated scenarios.
-In addition, our analysis revealed that substituting the transformation matrix with our neural network approach resulted in the most significant enhancement in the efficacy of the debiasing method.
-This observation underscores the pivotal role played by neural networks in maximizing the effectiveness of the debiasing techniques.
-We present a visualization of the results of our Ablation experiments in Table 5.
-This empirical evidence underscores the robustness and efficacy of our proposed approach in mitigating bias within natural language processing systems.
-By combining state-of-the-art debiasing techniques with advanced optimization strategies, we have unlocked a powerful methodological framework for enhancing the fairness and accuracy of language models.
-8 Discussion
-In this section, we summarise the answers to our research questions.
-8.1 RQ1
-We find that DeepSoftDebias outperforms state-of-the-art methods, and does so without negatively affecting downstream task performance.
-We make this conclusion after exhaustive testing on several models, and datasets and evaluating several metrics.
-8.2 RQ2
-We find that size and complexity do affect the ability of debiasing models. Specifically, we make the following observations about DeepSoftDebias:
-  • A single layer neural network can effectively de-bias embeddings with $\text{dim}\leq 1024$.
-  • A two-layer neural network can effectively debias embeddings with $\text{dim}\leq 2048$.
-  • A two-layer neural network with an increased layer size can effectively de-bias embeddings with $\text{dim}\leq 4096$.
-  • A three-layer neural network can effectively debias embeddings with $\text{dim}\leq 4450$.
-As a step for future work, we are curious to investigate scaling patterns to a further extent.
-8.3 RQ3
-While debiasing techniques in general can affect the downstream performance of models, we test DeepSoftDebias on multiple challenging downstream tasks and report that our proposed approach, to a large extent, does not negatively influence the performance of different downstream tasks.
-Remarkably, we see an improvement when using our debiased embeddings for some downstream tasks.
-8.4 RQ4
-We find that while DeepSoftDebias is effective at reducing bias across gender, race, and religion.
-We conclude this after testing on multiple embeddings, and multiple datasets and evaluating on multiple performance metrics.
-As a step for future work, we are curious to investigate whether our proposed approach works towards other forms of bias as well.
-9 Conclusion
-In this paper, we propose DeepSoftDebias, an approach that leverages neural networks to reduce bias in large language model embeddings.
-We perform an exhaustive series of tests using multiple performance metrics, state-of-the-art datasets, and downstream tasks to ensure that our debiasing technique is robust, efficient, and accurate.
-In the future, it would be interesting to see how this method translates to multilingual datasets since bias is language and culture-specific.
-We hope that this research paves the way for future endeavors that look to make large language models fair, ethical, and bias-free.
-10 Limitations
-While we do perform exhaustive analysis to test our proposed methodology, our study is monolingual and covers datasets only in English.
-Consequently, our downstream tasks are also tested only in English.
-Further, we were unable to test on API-based models at this time.
-11 Ethics Statement
-We understand that bias can be defined in various ways, and it’s not necessarily ideal for a language model to treat all users exactly the same without considering demographics.
-There are situations where certain topics require careful handling to avoid perpetuating harmful stereotypes against marginalized communities.
-Using specific bias metrics might suggest they encompass all negative social impacts across different groups, but we recognize that existing metrics may not capture all nuances in treatment across demographics.
-Therefore, any benchmark for bias needs to continually evolve to better understand and address these issues as they affect different communities.
-The definitions of morality and bias are shaped by cultural perspectives, resulting in diverse interpretations among individuals.
-Consequently, we do not claim that this work provides an objective or exhaustive measure of any of these concepts.
-Appendix
-This section provides supplementary material in the form of additional examples, implementation details, etc. to bolster the reader’s understanding of the concepts presented in this work.
-Appendix A MAC Scores of DeepSoftDebias
-Figures 3, 4, and 5 illustrate how DeepSoftDebias reduces bias in LLM embeddings.
-Figure 3: A visual representation of how DeepSoftDebias reduces gender bias in large language model embeddings.
-Figure 4: A visual representation of how DeepSoftDebias reduces racial bias in large language model embeddings.
-Figure 5: A visual representation of how DeepSoftDebias reduces religion bias in large language model embeddings.
-Appendix B Stereoset Scores of DeepSoftDebias
-Figures 6 and 8 provide an illustration of word vectors debiased using DeepSoftDebias and their stereoset scores and Crows Metric scores respectively.
-Figure 6: A visual representation of word vectors debiased using DeepSoftDebias and their stereotype scores across gender, race and religion respectively.
-Figure 7: A visual representation of word vectors debiased using DeepSoftDebias and their Crows Metric score across gender, race and religion respectively.
-Figure 8: A visual representation of word vectors debiased using DeepSoftDebias and their Crows Metric scores across gender, race and religion respectively.
-Appendix C Downstream Testing Results
-Figures 9 and 10 present an illustration of the results of various downstream tasks and their performance evaluation.
-Figure 9: An illustration of the results of downstream testing on NER. We compare the performance of biased and debaised embeddings in the directions of gender, race, and religion respectively.
-Figure 10: An illustration of results of downstream testing on sentiment analysis.
-We compare the performance of biased and debaised embeddings in the directions of gender, race, and religion respectively.
-Figure 11: An illustration analysis of number of layers in debiasing neural network vs.
-embedding dimension.
-We can see the varying performance of the 3 different sizes according to the embedding dimension of the LM it is used with.

### v5

@@
 We exhaustively evaluate this algorithm across a variety of state-of-the-art datasets, accuracy metrics, and challenging NLP tasks.
-We find that DeepSoftDebias outperforms the current state-of-the-art methods at reducing bias across gender, race, and religion.
+On a wide range of metrics, we find that DeepSoftDebias outperforms the current state-of-the-art methods at reducing bias across gender, race, and religion.
 1 Introduction
@@
 They act as the basis for these models to understand and subsequently, generate human-like language.
-However, it has been shown that these word embeddings themselves may reflect or exhibit bias Dev et al.
+However, it has been shown that these word embeddings may reflect or exhibit bias Dev et al.
 (2020); May et al.
@@
 (2014).
-This study prompted the NLP community to reevaluate the fundamental choices underlying our word representation models.
+This study prompted the NLP community to re-evaluate the fundamental choices underlying our word representation models.
 Specifically, we present DeepSoftBias: an algorithm that furthers the application of their methodology, by diverging from the conventional GloVe embeddings and delving into the word embeddings produced by the best-performing models on the Massive Text Embedding Benchmark (MTEB) Muennighoff et al.
@@
 To this end, our work answers the following research questions:
-  1. RQ1: Compared to traditional methods, does our proposed methodology attain better performance metrics when it comes to debiasing SOTA model embeddings?
-  2. RQ2: How do parameters of the model (size, complexity) interact with various SOTA debiasing techniques? What effect do they have on each other?
-  3. RQ3: To what extent do various SOTA debiasing techniques influence the performance of models on different downstream tasks?
+  1. RQ1: Compared to traditional methods, does our proposed methodology attain better performance metrics with respect to debiasing SOTA model embeddings?
+  2. RQ2: How does our proposed method interact with the varying parameters (size, complexity, embedding dimension) of embeddings obtained from different language models ?
+  3. RQ3: Can we validate that the debiased embeddings, which are a result of our proposed method, are on par with off-the-shelf embeddings on downstream tasks?
   4. RQ4: How does the type of bias (gender, race, religion) affect the effectiveness of the debiasing process?
 To answer the above questions, we make the following contributions through this research:
-Our Contributions We provide, to the best of our knowledge, the first comprehensive study of how various debiasing methods work on SoTA LLM word embeddings We present a novel methodology, DeepSoftDebias, for debiasing LLM word embeddings, which beats SoTA debiasing methods across multiple bias formats including gender, race, and religion.
+Our Contributions We provide, to the best of our knowledge, the first comprehensive study of how various debiasing methods work on SoTA large language model word embeddings.
+We present a novel methodology, DeepSoftDebias, for debiasing LLM word embeddings, which beats SoTA debiasing methods across multiple bias formats including gender, race, and religion.
 We perform an exhaustive quantitative analysis, establishing SoTA baselines and leveraging multiple evaluation metrics to provide a comparison against accessible SoTA baselines.
@@
 1.
-We find that DeepSoftDebias not only outperforms the state-of-the-art methods at reducing bias across gender, race, and religion but also does so while preserving the full information of the original embedding (which is an additional improvement on previous methods).
-Further, we find that model performance on challenging downstream tasks like NER and sentiment analysis remains largely unaffected when we test using our debiased embeddings.
+We find that DeepSoftDebias not only outperforms the state-of-the-art methods at reducing bias on most of the bias types but also does so while preserving the full information of the original embedding (which is an additional improvement on previous methods).
+Notably, our proposed methodology is also effective for debiasing embeddings from state space models, namely Mamba Gu and Dao (2024).
+Further, we find that model performance on challenging downstream tasks like the ones present in the GLUE Benchmark Wang et al.
+(2018) remains largely unaffected when we test using our debiased embeddings.
+We also make all of our code available at https://github.com/aishik-rakshit/DeepSoftDebias
 Figure 1: A step-by-step visualization of the pipeline for DeepSoftDebias.
@@
 The INLP methodology debiases representations by iteratively projecting them into a linear classifier’s null space.
-This technique is particularly effective for handling intersectional groups, which are defined by combinations of sensitive attributes².
-INLP seeks to learn a hidden representation that is independent of the protected attributes.
 This approach is beneficial in scenarios where an attempt to make a model fairer towards some group results in increased unfairness towards another group.
-Therefore, INLP emerges as a robust and effective strategy for mitigating bias in language models, promoting fairness across multiple protected attributes.
 Self-Debias
@@
 (2021) is a novel approach to mitigating bias in language models.
-The methodology, first coined by Schick et al.
-(2021), is based on the concept of self-diagnosis.
+The methodology is based on the concept of self-diagnosis.
 In this approach, pretrained language models recognize their undesirable biases and the toxicity of the content they produce.
@@
 While it does not completely eliminate the issue of language models generating biased text, it is an important step in this direction.
-The self-debiasing approach demonstrates the potential of language models to self-regulate and reduce their inherent biases.
 Sentence Debias
@@
 This process effectively debiases the sentence representations.
-SDB is particularly useful for mitigating biases related to gender, race, and religion.
-It offers a comprehensive comparison between models that adjust weights for debiasing and those employing test-time surgical interventions.
-The SDB method signifies a significant advancement in debiasing strategies, promoting a more equitable representation in language models.
+,offering a comprehensive comparison between models that adjust weights for debiasing and those employing test-time surgical interventions.
 Counterfactual Data Augumentation
@@
 This technique is part of a broader set of debiasing techniques that also includes Dropout, Self-Debias, SentenceDebias, and Iterative Nullspace Projection.
-CDA has been applied to various language models, including BERT, with the goal of diminishing stereotypical biases while maintaining the model’s performance on downstream tasks.
-However, it’s important to note that while CDA has the potential to improve the fairness of NLP models, it may not be effective in eliminating all biases and may even introduce new biases or errors in the model³.
 FineDeb
@@
 In the second phase, the debiased model is fine-tuned on the language modeling objective.
-This methodology is effective for demographics with multiple classes.
-The FineDeb approach demonstrates its effectiveness through extensive experiments and comparisons with state-of-the-art techniques.
-It offers stronger debiasing in comparison to other methods, which often result in models as biased as the original language model.
-Thus, FineDeb emerges as a robust and effective framework for mitigating bias in language models.
+Though FineDeb emerges as a robust and effective framework for mitigating bias in language models, it is not open sourced at this time.
+One of our goals with DeepSoftDebias is to offer an at least equally performative alternative that is open sourced.
 3 Data
@@
 (2018) corpus, a collection of Reddit posts and comments by both native and non-native English speakers, featuring approximately 56 million sentences.
-This dataset serves as the foundation for training word embeddings, aiming to capture the nuanced and inherently biased linguistic patterns present in social media discourse.
+This dataset serves as our foundation for training word embeddings, aiming to capture the nuanced and inherently biased linguistic patterns present in social media discourse.
 In our study, we employ the Reddit L2 corpus as the source for our initial Word2Vec Mikolov et al.
@@
 The primary objective of this approach is to ensure a consistent set of words across all our LLMs.
-This consistency allows each of our LLMs to be tested on the same set of words.
 StereoSet
@@
 (2019), RoBERTa Liu et al.
-(2019), and XLNet Yang et al.
+(2019a), and XLNet Yang et al.
 (2020).
@@
 In this section, we delve into the domain of debiasing word embeddings, presenting both an established and a newly proposed methodology for mitigating biases in word vector representations.
-These biases span across gender, racial, and religious lines and are encoded inadvertently within language models.
 4.1 Bias Identification and Data Structure
@@
 To enhance performance, we propose DeepSoftDebias.
-In this approach, we replace the transformation matrix with a neural network, leveraging its capability to represent a sequence of transformation matrices.
+In this approach, we replace the transformation matrix with a neural network made up of residual blocks He et al.
+(2015), leveraging its capability to represent a sequence of transformation matrices.
 This adaptation enables the algorithm to handle more complex functions mapping between input and output embeddings.
@@
 We demonstrate the effectiveness of DeepSoftDebias on various datasets and tasks, and show that it outperforms the state-of-the-art methods in terms of accuracy and efficiency.
-The reason for the need of a fixed Biasspace is that we adopt the methodology proposed by Bolukbasi et al.
+The reason for the need for a fixed BiasSpace is that we adopt the methodology proposed by Bolukbasi et al.
+(2016).
 for the derivation of the bias subspace.
+The Fixed BiasSpace is crucial for mitigating bias in word embeddings by providing a fixed subspace representing the direction of bias.
+The neural network is trained to make embeddings orthogonal to this subspace, reducing bias while maintaining other semantic relationships.
+This orthogonality minimizes the projection of words onto the bias subspace.
 The process of creating the BiasSpace commences with the identification of word vectors representing opposing concepts, such as ‘he’ versus ‘she’, or ‘man’ versus ‘woman’.
@@
 2 presents a visualization of our approach to downstream testing.
+Our methodology also extends to multi-dimensional bias representations across race, religion, and other social dimensions.
+For each group set, we compute a neutral reference point by calculating the mean vector of representative terms.
+For instance, in racial bias, we derive a neutral vector from terms like "black," "caucasian," and "asian," while for religious bias, we use terms representing different faith traditions such as "judaism," "christianity," and "islam." We then compute bias-specific axes by subtracting this neutral reference vector from each group’s representative vector, yielding vectors that exclusively capture bias components.
+By taking the vector differences—such as $v_{a}sian$ $v_{n}eutral$, $v_{w}hite$ $v_{n}eutral$, and $v_{b}lack$ $v_{n}eutral$—we create directional axes that quantify bias in the embedding space.
+These bias vectors are concatenated to form a comprehensive bias subspace matrix, which plays a pivotal role in the training of our neural network.
+Specifically, we ensure that the output of the word embeddings, upon being processed through the neural network, is orthogonal to this bias subspace.
+This approach allows us to systematically identify, measure, and mitigate bias across multiple social dimensions, moving beyond binary conceptualizations to a more nuanced representation of semantic prejudices.
+Figure 2 presents a visualization of our approach to downstream testing, illustrating how our method can be applied to diverse contextual scenarios.
+We also provide a schematic of an example of the transformation Neural Network in Figure 12
 Input: Biased word embeddings ($\text{emb}_{\text{biased}}$), Bias Subspace (BiasSpace), Neutral word embeddings ($\text{emb}_{\text{neutral}}$)
@@
 We observe in 11 Smaller models, such as bge-small Xiao et al.
-(2023) and DeBERTa-v3-base He et al.
-(2023) or DeBERTa-v3-large, can be debiased effectively by a single-layer neural network.
+(2023), DeBERTa-v3-base He et al.
+(2023) or DeBERTa-v3-large, GPT2Radford et al.
+(2019) and RobertaLiu et al.
+(2019b) can be debiased effectively by a single-layer neural network.
 Larger models, such as Llama-2 Touvron et al.
-(2023), Alpaca Taori et al.
-(2023) and Yi-6b 01.ai (2024) need a more complex debiasing neural network.
-For embeddings with embedding length of around 2000, a two-layer neural network is sufficient, while for larger embedding dimensions, a three-layer neural network is required to achieve good debiasing results.
+(2023), Llama-3 AI@Meta (2024), Alpaca Taori et al.
+(2023) and Yi-6b 01.ai (2024), GemmaTeam et al.
+(2024), QwenBai et al.
+(2023) and MambaGu and Dao (2024) need a more complex debiasing neural network.
+For embeddings with an embedding length of around 2000, a two-layer neural network is sufficient, while for larger embedding dimensions, a three-layer neural network is required to achieve good debiasing results.
 In addressing the second research question, we delve into the intricacies of neural network complexity necessary for debiasing embeddings of varying sizes.
@@
 In this section, we provide an extensive analysis of our proposed methodology, complete with a comprehensive evaluation against multiple metrics, tasks, and datasets.
-We provide the results of additional downstream testing and ablation experiments in Appendix D and Appendix F, respectively.
+We provide the results of additional downstream testing and ablation experiments in Appendix D and Appendix H, respectively.
 We also provide our hypothesis of why there is a variation in bias across LLMs in Appendix E.
-Model Variant Topic BMAC NSMAC SS CMS CSS CAS
-Yi Yi-6B Gender 0.148 0.964 55.372 49.620 58.970 37.250
-Alpaca Alpaca-7B 0.612 0.816 53.306 48.850 57.690 37.250
-BAAI bge-base-en-v1.5 0.471 0.997 50.000 48.090 42.310 58.820
-BAAI bge-large-en-v1.5 0.404 0.983 49.174 50.380 50.640 51.960
-Zephyr Zephyr-7B-beta 0.393 0.981 52.893 46.950 59.620 29.410
-Mistral e5-mistral-7b-instruct 0.343 0.971 52.893 48.090 55.770 38.240
-Llama 2 Llama-2-7b-hf 0.182 0.964 48.347 44.660 57.690 26.470
-Salesforce SFR-Embedding-Mistral 0.343 0.971 50.000 45.420 50.000 40.200
-Falcon falcon-7b 0.011 0.964 51.240 48.850 60.900 32.350
-Gemma Gemma-2b 0.058 0.971 47.107 48.470 57.69 36.27
-Gemma Gemma-7b 0.553 0.976 49.173 51.53 63.46 35.29
-GritLM GritLM-7B 0.379 0.999 51.239 48.470 57.05 37.25
-mxbai mxbai-embed-large-v1 0.467 0.994 51.652 55.34 60.9 49.02
-Yi Yi-6B Race 0.111 0.964 46.209 64.150 66.170 53.660
-Alpaca Alpaca-7B 0.655 0.938 52.357 41.280 41.540 46.340
-BAAI bge-base-en-v1.5 0.496 0.992 49.590 44.770 46.250 36.590
-BAAI bge-large-en-v1.5 0.404 0.990 50.922 40.890 40.690 51.220
-Zephyr Zephyr-7B-beta 0.419 0.992 49.283 42.250 41.330 60.980
-Mistral e5-mistral-7b-instruct 0.380 0.999 50.922 52.520 52.680 60.980
-Llama 2 Llama-2-7b-hf 0.175 0.990 50.410 45.930 46.680 46.340
-Salesforce SFR-Embedding-Mistral 0.381 0.994 51.639 49.030 50.750 39.020
-Falcon falcon-7b 0.010 0.985 50.922 46.710 46.900 53.660
-Yi Yi-6B Religion 0.147 0.984 52.564 47.620 48.480 33.330
-Alpaca Alpaca-7B 0.676 0.823 51.282 80.000 82.830 33.330
-BAAI bge-base-en-v1.5 0.497 0.990 46.154 59.050 61.620 16.670
-BAAI bge-large-en-v1.5 0.406 0.985 51.282 60.000 61.620 33.330
-Zephyr Zephyr-7B-beta 0.465 0.996 51.282 48.570 50.510 16.670
-Mistral e5-mistral-7b-instruct 0.436 0.985 52.564 52.380 51.520 66.670
-Llama 2 Llama-2-7b-hf 0.202 1.003 44.872 64.760 66.670 33.330
-Salesforce SFR-Embedding-Mistral 0.437 0.988 51.282 40.950 39.390 66.670
-Falcon falcon-7b 0.009 0.998 48.718 50.480 51.520 33.330
-Table 1: Quantitative analysis for DeepSoftDebias using BiasedMAC (BMAC), New SoftMAC (NSMAC), StereotypeScore (SS), Crows Metric Score (CMS), Crows Stereotype Score (CSS), Crows Antistereotype Score (CAS).
-The best performance is highlighted in bold.
+Bias Type Model Name Biased MAC Soft-Debiased MAC DSD MAC Self-Debias MAC INLP MAC Sent-Debias MAC
+Gender Bert-Large-Uncased 0.343 0.734 0.917 - 0.720 0.453
+Roberta-Base 0.025 0.195 0.890 - 0.624 0.325
+Gemma-7b 0.540 0.160 0.908 - - -
+Llama 3 8b 0.254 - 0.939 - - -
+SFR Embedding 2_R 0.388 - 0.971 - - -
+Mamba 2.8b 0.101 - 0.919 - - -
+Race Bert-Large-Uncased 0.440 0.798 0.997 - 0.353 0.450
+Roberta-Base 0.025 0.205 0.863 - 0.422 0.524
+GPT2-xl 0.497 0.910 0.945 0.378 0.494 0.468
+Gemma-7b 0.546 0.111 0.951 - - -
+Llama 3 8b 0.252 - 0.937 - - -
+SFR Embedding 2_R 0.440 - 0.994 - - -
+Mamba 2.8b 0.109 - 0.958 - - -
+Religion Bert-Large-Uncased 0.455 0.745 0.971 - 0.326 0.482
+Roberta-Base 0.025 0.209 0.936 - 0.423 0.425
+GPT2-xl 0.477 0.927 0.977 0.343 0.549 0.561
+Gemma-7b 0.372 - 0.907 - - -
+Llama 3 8b 0.264 - 0.951 - - -
+SFR Embedding 2_R 0.455 - 0.988 - - -
+Mamba 2.8b 0.087 - 0.945 - - -
+Table 1: Debiasing Results showcasing the MAC scores for the 5 different Debiasing Methods
 6.1 Mean Average Cosine Similarity
@@
 We provide visual representations of the efficiency of DeepSoftDebias at removing gender bias, racial bias, and religion bias in Appendix B.
-In our research, we have chosen to utilize Mean Average Cosine Similarity (MAC) as our primary metric for assessing bias in word embeddings.
+In this work, we chose to utilize Mean Average Cosine Similarity (MAC) as our primary metric for assessing bias in word embeddings.
 This decision is informed by the work of Manzini et al.
 (2019), who posit that MAC can be viewed as an extension of the Word Embedding Association Test (WEAT), specifically adapted for a multiclass setting.
-The MAC and WEAT serve distinct, yet complementary purposes.
 While WEAT is designed to focus on specific associations between word vectors and predefined concepts (such as gender or race), MAC provides a broader perspective by measuring overall similarity patterns across different groups.
@@
 Table 2 presents DeepSoftDebias’s results alongside existing approaches on the Stereoset dataset.
-Notably, DeepSoftDebias achieves the lowest $SS$ across all social groups, demonstrating its effectiveness in bias reduction.
+Notably, DeepSoftDebias achieves the lowest $SS$ across most social groups, demonstrating its effectiveness in bias reduction.
 Particularly impressive is DeepSoftDebias’s performance in the gender and race categories, where it significantly outperforms existing methods.
-For instance, with the SFR-Embedding-Mistral Jiang et al.
-(2023) model, DeepSoftDebias achieves an $SS$ of 50 for gender and 50.409 for race using the Llama-2-7b model.
-Additionally, DeepSoftDebias attains a score of 51.282 for the Zephyr-7b-beta Tunstall et al.
-(2023) or 48.717 for Alpaca-7b Taori et al.
-(2023).
-We present these score in 2 an illustration of these scores in Fig.
+For instance, with the Mamba 2.8b Jiang et al.
+(2023) model, DeepSoftDebias achieves an $SS$ of 50 for gender and with the Llama 3 8b model it acheives a SS of 49.8 for race.
+We present these scores in Table 2 and an illustration of these scores in Fig.
 6.
@@
 Sentence Debias 59.37 57.78 58.73
-DeepSoftDebias 50.00 50.41 51.28
+DeepSoftDebias 50.00 49.8 52.56
 Table 2: StereoSet evaluation. Closer to 50 is better for SS. The best performance is highlighted in bold while the next best is underlined).
 6.3 Crows-Pairs Dataset
-Our study evaluates social bias in natural language generation tasks using the CrowS Pairs dataset, comprising pairs of sentences differing in their degree of bias.
+Our study also evaluates social bias in natural language generation tasks using the CrowS Pairs dataset, comprising pairs of sentences differing in their degree of bias.
 By ranking these sentences according to bias level, we quantify the effectiveness of various methods in reducing bias in word embeddings.
@@
 We then see whether the context vector is closer to the more (anti-)stereotypical sentence or the less (anti-)stereotypical sentence.
-If it is closer to the more (anti–)stereotypical sentence,, then we state that it agrees with the (anti–)stereotype, i.e., the label direction.
+If it is closer to the more (anti-)stereotypical sentence2901, then we state that it agrees with the (anti-)stereotype, i.e., the label direction.
 Particularly noteworthy is DeepSoftDebias’s superior performance in the gender and religion categories.
-For instance, with the Yi-6B model, DeepSoftDebias achieves a score of 49.62 for gender and 50.48 for religion with the falcon-7b model.
-Similarly, using the SFR-Embedding-Mistral model, DeepSoftDebias achieves a score of 49.03 for race biasing the SFR-Embedding-Mistral model.
+For instance, with the mamba-1.4b model, DeepSoftDebias achieves a score of 50.38 for gender and 50.48 for religion with the mamba-2.8b model.
+Similarly, using the bge-base-en v1.5 model, DeepSoftDebias achieves a score of 50.19 for debiasing for the bias-type of race.
 These results underscore the effectiveness of DeepSoftDebias in mitigating social bias in word embeddings.
-We present these score in 3 and depict the variation of these scores in Fig.
+We present these scores in 3 and 5 and depict the variation of these scores in Fig.
 8.
-We also report the CSS and CAS score which refer to the CrossNER Stereotype score, i.e., the number of times the model agrees with the more stereotypes statement when the label direction is stereotype, and the CrossNER Anti-stereotype score, which refers to the number of times the model agrees with the more anti-stereotyped statement when the label direction was anti-sterotype.
 Crows Pairs Metric Score (CMS)
@@
 Sentence Debias 52.29 62.72 63.81
-DeepSoftDebias 50.38 49.07 50.48
+DeepSoftDebias 50.38 50.19 50.48
 Table 3: Crows Pairs evaluation. Metric score for every demographic. Closer to 50 is better for the metric (best; next best).
+6.4 Downstream Testing
+For our downstream evaluation, we utilized the GLUE benchmark Wang et al.
+(2018).
+We present the performance differentials between the original word embeddings and their debiased counterparts, processed through various debiasing methods focusing on the three categories of GLUE tasks: single-sentence tasks, sentence-pair tasks, and inference-based tasks.
+For the single-sentence task, specifically the Stanford Sentiment Treebank (SST), we report accuracy as the primary metric in Table 6.
+In the case of sentence-pair tasks, exemplified by the Microsoft Research Paraphrase Corpus (MRPC), we use the F1 score as our performance indicator in Table 8.
+For the inference-based tasks, which include QNLI (Question-answering Natural Language Inference), WNLI (Winograd Natural Language Inference), RTE (Recognizing Textual Entailment), and MNLI (Multi-Genre Natural Language Inference), we report the average delta of the F1 scores across these four tasks in Table 7.
+For inference-based tasks we see a average gain in performance of 0.017 F1 Score for gender, followed by 0.027 for religion, and a net average performance delta of 0 for race.
+For sentence pair tasks, we see a performance delta of 0.047 for gender, 0.035 for race and -0.037 F1 score for religion.
+Our results indicate that the DeepSoft Debias method yields performance comparable to that of the original biased word embeddings, with variations in performance ranging from approximately 2% to 4% across the different metrics.
+We hypothesize that this outcome is attributable to the varying degrees of bias present in the datasets, which the models have become less susceptible to following the debiasing process.
+We have provided a detailed score $\Delta$s for all the models we have applied out DeepSoftDebias debiasing method on with the GLUE Tasks in Tables 9, 10, 11, 12, 13, 14 and 16 in the Appendix.
 7 Discussion
@@
 We find that size and complexity do affect the ability of debiasing models. Specifically, we make the following observations about DeepSoftDebias:
-  • A single layer neural network can effectively de-bias embeddings with $\text{dim}\leq 1024$.
-  • A two-layer neural network can effectively debias embeddings with $\text{dim}\leq 2048$.
-  • A two-layer neural network with an increased layer size can effectively de-bias embeddings with $\text{dim}\leq 4096$.
-  • A three-layer neural network can effectively debias embeddings with $\text{dim}\leq 4450$.
-As a step for future work, we are curious to investigate scaling patterns to a further extent.A visualization of this is provided in Fig 11
+  • A single residual block(RB) neural network can effectively de-bias embeddings with $\text{dim}\leq 1024$.
+  • A two RB neural network can effectively debias embeddings with $\text{dim}\leq 2048$.
+  • A three RB neural network with an increased layer size can effectively de-bias embeddings with $\text{dim}\leq 4450$.
+With respect to future work, we are curious to investigate scaling patterns to a further extent.
+A visualization of this is provided in Fig 11.
+We offer a further detailed discussion on hyperparameter tuning the DeepSoftDebias Method in Question 5 of our FAQ section.
+2
 RQ3
@@
 As a step for future work, we are curious to investigate whether our proposed approach works towards other forms of bias as well.
+We offer a further detailed quantitative discussion of the bias types it works best with, in Question 6 of our FAQ Section.
 8 Conclusion
@@
 These models could not be accommodated within the GPU VRAM, even after applying quantization to 8 bits.
-Consequently, the largest model that we were able to test was the Gemma-7B model.
+Consequently, the largest model that we were able to test was the Gemma-2-9b model.
 10 Ethics Statement
@@
 Race Black "slave", "musician", "runner", "criminal", "homeless"
-Asian "manager", "executive", "redneck", "hillbilly", "leader", "farmer"
-Caucasian "doctor", "engineer", "laborer", "teacher"
+Caucasian "manager", "executive", "redneck", "hillbilly", "leader", "farmer"
+Asian "doctor", "engineer", "laborer", "teacher"
 Religion Jew "greedy", "cheap", "hairy", "liberal"
@@
 Figures 6 and 8 provide an illustration of word vectors debiased using DeepSoftDebias and their stereoset scores and Crows Metric scores respectively.
+We present the MAC scores, Stereotype Scores, Crows-Pairs Metric Scores in Table 5
+Bias Type Model Name Biased MAC Soft-Debiased MAC DSD MAC Soft-Debiased SS DSD SS Soft-Debiased CMS DSD CMS
+gender gte-Qwen2-7B-instruct 0.469 - 0.95 25.21 45.04 32.82 41.22
+bge-base-en-v1.5 0.447 0.88 0.966 47.93 43.39 43.13 45.04
+bge-large-en-v1.5 0.408 0.881 0.921 49.59 48.35 41.98 50.76
+roberta-base 0.025 0.179 0.922 52.06 50.41 53.44 51.53
+bert-base-uncased 0.243 0.408 0.921 50.41 52.07 40.84 45.8
+bert-large-uncased 0.324 0.734 0.917 52.89 48.76 50 48.85
+gemma-2-2b 0.46 - 0.894 25.62 49.59 26.72 37.4
+gemma-2-9b 0.502 - 0.915 25.21 48.76 30.15 46.95
+gemma-2b 0.056 - 0.847 19.42 47.11 28.24 41.98
+gemma-7b 0.54 0.16 0.908 45.45 48.35 45.04 45.8
+GritLM-7B 0.38 - 0.905 24.38 49.59 25.95 39.31
+Meta-Llama-3-8B 0.254 - 0.939 19.83 50 23.66 49.24
+gpt2-xl 0.497 0.91 0.945 49.17 48.35 46.18 49.62
+SFR-Embedding-2_R 0.388 - 0.906 26.03 51.24 30.53 40.08
+mamba-1.4b-hf 0.342 - 0.935 43.39 46.28 40.08 50.38
+mamba-2.8b-hf 0.101 - 0.919 21.9 50 18.32 51.91
+race gte-Qwen2-7B-instruct 0.426 - 0.971 26.02 51.23 16.86 63.18
+bge-base-en-v1.5 0.467 0.903 0.987 51.84 51.02 60.27 50.19
+bge-large-en-v1.5 0.424 0.938 0.99 51.33 51.43 39.34 50.78
+roberta-base 0.025 0.205 0.863 48.77 47.44 30.62 47.04
+bert-base-uncased 0.245 0.448 0.974 52.25 51.43 29.07 46.07
+bert-large-uncased 0.354 0.798 0.997 49.28 49.69 61.43 44.96
+gemma-2-2b 0.44 - 0.966 29.3 47.95 41.67 56.2
+gemma-2b 0.046 - 0.838 21.31 50.31 12.98 51.16
+gemma-7b 0.546 0.111 0.951 47.75 46.72 53.1 43.22
+GritLM-7B 0.417 - 0.971 21.41 51.23 15.5 46.12
+Meta-Llama-3-8B 0.252 - 0.937 22.23 49.8 27.91 43.99
+gpt2 0.01 - 0.822 20.49 48.77 12.79 38.18
+gpt2-xl 0.477 0.927 0.977 49.59 51.02 50.39 55.62
+SFR-Embedding-2_R 0.44 - 0.99 29.2 51.74 32.75 59.88
+mamba-1.4b-hf 0.356 - 0.994 45.7 49.39 56.98 47.29
+mamba-2.8b-hf 0.109 - 0.958 21.11 51.33 19.19 52.52
+religion gte-Qwen2-7B-instruct 0.442 - 0.975 21.79 57.69 19.05 70.48
+bge-base-en-v1.5 0.474 0.897 0.987 50 47.44 49.52 61.9
+bge-large-en-v1.5 0.412 0.895 0.94 53.85 55.13 60 56.19
+roberta-base 0.025 0.209 0.936 51.28 56.41 36.19 34.29
+bert-base-uncased 0.276 0.477 0.991 56.41 46.15 32.38 68.57
+bert-large-uncased 0.312 0.745 0.971 52.56 48.72 42.86 60.95
+gemma-2-2b 0.523 - 0.957 20.51 52.56 29.52 44.76
+gemma-2b 0.095 - 0.933 24.36 55.13 12.38 58.1
+gemma-7b 0.372 - 0.907 43.59 52.56 40.95 55.48
+GritLM-7B 0.485 - 0.963 17.95 56.41 25.71 55.24
+Meta-Llama-3-8B 0.264 - 0.951 20.51 58.97 20 52.43
+gpt2 0.018 - 0.953 12.82 52.56 22.86 63.81
+SFR-Embedding-2_R 0.455 - 0.959 34.62 55.13 41.9 54.29
+mamba-1.4b-hf 0.267 - 0.97 55.13 55.13 41.9 49.52
+mamba-2.8b-hf 0.087 - 0.945 19.23 46.15 27.62 50.48
+Table 5: Debiasing Results on LLMs: MAC, CMS and SS on all the models we have tested our method on
 Figure 6: A visual representation of word vectors debiased using DeepSoftDebias and their stereotype scores across gender, race and religion respectively.
@@
 Appendix D Downstream Testing Results
-Model Variant Baseline Debiased Text Class. Acc. DSB Debiased Text Class. Acc. Baseline Debiased NER Macro F1 Avg. DSB Debiased NER Macro F1 Avg.
-Gemma gemma-2b 0.7655 0.7964 0.469 0.484
-BAAI bge-base-en-v1.5 0.8296 0.822 0.458 0.421
-Mistral SFR-Embedding-Mistral 0.8297 0.821 0.404 0.428
-Gemma gemma-7b 0.516 0.8032 0.198 0.475
-Zephyr zephyr-7b-beta 0.7997 0.81 0.403 0.429
-mxbai mxbai-embed-large-v1 0.8366 0.7903 0.461 0.455
-Table 5: Downstream testing results comparison with embeddings debiased using DeepSoftDebias and the baseline SoftDebais Method.
-The first two columns represent results for downstream performance on sentiment analysis.
-The second two columns represent results for downstream performance on NER.
 In our research, we primarily focus on the debiasing of word embeddings derived from Language Learning Models (LLMs).
@@
 This approach allows us to provide a more comprehensive understanding of the implications and potential benefits of debiasing word embeddings.
-D.1 Sentiment Classification
+D.1 Text Classification
 In our study, we employ downstream testing to assess the utility of embeddings debiased using DeepSoftDebias across two key natural language processing tasks: text classification and named entity recognition (NER).
-Utilizing the IMDB Sentiment Classification dataset İlhan Tarımer et al.
-(2019) and Stanford TreeBank Dataset for text classification, featuring labeled movie reviews as positive or negative, we compute the average sentence vectors using both original and debiased embeddings.
+Utilizing the Stanford TreeBank DatasetSocher et al.
+(2013) for text classification.
 Training XGBoost Chen and Guestrin (2016) classifiers on these vectors, we compare their accuracy on the test set, recognizing accuracy as a straightforward metric for binary classification tasks like sentiment analysis.
-Notably, our results reveal a performance improvement when debiasing in the gender and religion directions, whereas a slight decrease in performance is observed in the case of race debiasing.
-We provide these results in Table:6 for IMDB Sentiment classification and Table:7 for Stanford Sentiment Treebank.
+Notably, our results reveal a slight performance improvement when debiasing in the gender and religion directions inmost cases, whereas a slight decrease in performance is observed in the case of race debiasing in mostr cases.
+We provide these results in Table:9 for Stanford Sentiment Treebank.
 A visual representation of these results in Fig.
 9.
-D.2 Named Entity Recognition (NER)
-In our research, we examine the performance of debiased embeddings in the domain of named entity recognition (NER) using the Reuters subset of the CrossNER Liu et al.
-(2020) dataset.
-This dataset comprises news domain sentences annotated with four entity types: person, location, organization, and product.
-Employing a simple BiLSTM model, we input padded arrays of embeddings for each sentence and trained the model on the dataset.
-We evaluate the models’ performance on the test set using the macro-averaged F1-score, a metric that balances precision and recall, crucial for accurate entity identification and classification.
-To mitigate potential bias towards more frequent entity types, we adopt macro-averaging, allotting equal importance to each entity type.
-Remarkably, our findings indicate a slight performance boost when using debiased embeddings in all three directions compared to biased embeddings.
-We provide these results in Table:6and a visual representation of these results in Fig.
-10.
-D.3 Semantic Textual Similarity
+D.2 Semantic Textual Similarity
 In our research, we evaluate the performance of debiased embeddings for the Semantic Textual Similarity (STS) task using the STS-B dataset.
@@
 This methodology allows us to develop and evaluate models’ ability to understand nuanced semantic relationships in text effectively.
-We provide our results in this task in Table:7
+We provide our results in this task in Table:10
+Bias Type Model Name $\Delta$ Soft-Debias $\Delta$ Self-Debias $\Delta$ INLP $\Delta$ Sent-Debias $\Delta$ DSD
+Gender Bert-Large-Uncased 1.17 - 0.453 0.067 -0.831
+Roberta-Base 2.75 - 0.2 0.252 0.527
+GPT2 -7.958 -3.133 -0.616 -0.831 -6.554
+Gemma-7b -11.352 - - - -2.633
+LLama 2 7b -1.229 - - - 0.059
+SFR Embedding 2_R 0.995 - - - 0.059
+Mamba 2.8b -4.623 - - - -0.995
+Race Bert-Large-Uncased 0.995 - 0.586 -0.03 -1.01
+Roberta-Base -0.293 - -0.171 0.104 0.761
+GPT2 -5.5 -2.992 -0.23 -1.01 -4.389
+Gemma-7b -10.474 - - - -1.697
+LLama 2 7b 0.995 - - - 0.41
+SFR Embedding 2_R 0.936 - - - 1.112
+Mamba 2.8b -5.734 - - - 0.059
+Religion Bert-Large-Uncased 1.053 - 0.483 -0.111 -1.514
+Roberta-Base 1.638 - 0.23 -0.245 1.989
+GPT2 0.585 -3.133 -0.29 -1.514 2.575
+Gemma-7b -13.517 - - - -2.867
+LLama 2 7b -1.112 - - - 0.293
+SFR Embedding 2_R 1.17 - - - 1.58
+Mamba 2.8b -4.096 - - - -1.814
+Table 6: Downstream Result: $\Delta$ of the Accuracy (out of 100) between downstream testing on the SST Dataset using the biased embedding and the Debiased embeddings using various debiasing methods
+Bias Type Model Name $\Delta$ Soft-Debias $\Delta$ Self-Debias $\Delta$ INLP $\Delta$ Sent-Debias $\Delta$ DSD
+Gender Bert-Large-Uncased -0.024 - 0.032 -0.113 -0.017
+Gemma-7b 0.024 - - - 0.008
+GPT2 0.005 0.0003 -0.005 0.031 -0.12
+Llama 3 8b -0.006 - - - 0.035
+Mamba 2.8b -0.018 - - - -0.007
+Roberta-Base 0.045 - 0.032 -0.113 0.045
+SFR Embedding 2_R 0.009 - - - 0.197
+Race Bert-Large-Uncased 0.04 - -0.001 0.007 -0.016
+Gemma-7b 0.015 - - - 0.001
+GPT2 0.04 0.036 -0.068 0.062 0.002
+Llama 3 8b -0.007 - - - 0.07
+Mamba 2.8b -0.009 - - - 0.009
+Roberta-Base 0.001 - -0.001 0.007 0.009
+SFR Embedding 2_R -0.192 - - - -0.219
+Religion Bert-Large-Uncased -0.012 - -0.052 -0.04 0.024
+Gemma-7b -0.053 - - - -0.06
+Llama 3 8b 0.002 - - - 0.009
+Mamba 2.8b 0.012 - - - 0.048
+Roberta-Base 0.011 - -0.052 -0.04 -0.021
+SFR Embedding 2_R -0.006 - - - 0.007
+Table 7: Downstream testing: Average $\Delta$ of the F1 Score between downstream testing on the the 4 Inference type tasks in GLUE (QNLI,WNLI,RTE,MNLI) using the biased embedding and the Debiased embeddings using various debiasing methods
+Bias Type Model Name $\Delta$ Soft-Debias $\Delta$ Self-Debias $\Delta$ INLP $\Delta$ Sent-Debias $\Delta$ DSD
+Gender Bert-Large-Uncased -0.162 -0.011 0.002 -0.146
+Gemma-7b 0.005 -0.007
+GPT2 0.166 0 -0.005 -0.004 -0.005
+Llama 3 8b -0.007 -0.16
+Mamba 2.8b 0.078 0
+Roberta-Base 0.014 0.014 -0.001 0.152
+SFR Embedding 2_R 0.146 0.144
+Race Bert-Large-Uncased 0.154 -0.003 0.014 -0.007
+Gemma-7b 0.238 0.245
+GPT2 -0.018 -0.002 -0.002 0.005 -0.001
+Llama 3 8b 0.153 -0.242
+Mamba 2.8b -0.164 -0.005
+Roberta-Base -0.014 0 -0.01 -0.021
+SFR Embedding 2_R 0 0.813
+Religion Bert-Large-Uncased -0.175 -0.008 0.005 -0.42
+Gemma-7b -0.14 0.015
+Llama 3 8b 0.018 0.003
+Mamba 2.8b 0 -0.006
+Roberta-Base -0.159 0 -0.01 -0.006
+SFR Embedding 2_R 0.241 0.235
+Table 8: Downstream testing: $\Delta$ of the F1 Score between downstream testing on MRPC Dataset using the Biased embedding and the Debiased embeddings using various debiasing methods
+Bias Type Model Name $\Delta$ Soft-Debias $\Delta$ DSD $\Delta$ Self-Debias $\Delta$ INLP $\Delta$ Sent-Debias
+gender Alibaba-NLP/gte-Qwen2-7B-instruct 0.351 1.463 - - -
+BAAI/bge-base-en-v1.5 -0.176 0.351 - - -
+BAAI/bge-large-en-v1.5 -1.872 -0.644 - - -
+FacebookAI/roberta-base 2.75 0.527 - 0.2 0.252
+google-bert/bert-base-uncased 0.468 -0.41 - 0.453 0.067
+google-bert/bert-large-uncased 1.17 1.755 - - -
+google/gemma-2-2b -6.963 -1.697 - - -
+google/gemma-2b -1.17 -0.117 - - -
+google/gemma-7b -11.352 -2.633 - - -
+GritLM/GritLM-7B -1.931 -0.995 - - -
+openai-community/gpt2-xl 0.585 2.575 -4.514 - -
+Salesforce/SFR-Embedding-2_R 0.995 0.059 - - -
+state-spaces/mamba-1.4b-hf -3.862 2.633 - - -
+state-spaces/mamba-2.8b-hf -4.623 -0.995 - - -
+race Alibaba-NLP/gte-Qwen2-7B-instruct 0.585 1.287 - - -
+BAAI/bge-base-en-v1.5 -0.234 0.234 - - -
+BAAI/bge-large-en-v1.5 -0.878 -1.872 - - -
+FacebookAI/roberta-base -0.293 0.761 - -0.171 0.104
+google-bert/bert-base-uncased 0.527 -0.351 - 0.586 -0.03
+google-bert/bert-large-uncased 0.995 2.165 - - -
+google/gemma-2-2b -7.139 -1.872 - - -
+google/gemma-2b -0.468 0.468 - - -
+google/gemma-7b -10.474 -1.697 - - -
+GritLM/GritLM-7B -0.702 -2.224 - - -
+openai-community/gpt2 -7.958 -6.554 -2.992 - -
+openai-community/gpt2-xl 0.819 1.463 -4.417 - -
+Salesforce/SFR-Embedding-2_R 0.936 1.112 - - -
+state-spaces/mamba-1.4b-hf -4.74 1.346 - - -
+state-spaces/mamba-2.8b-hf -5.734 0.059 - - -
+religion Alibaba-NLP/gte-Qwen2-7B-instruct -0.234 1.58 - - -
+BAAI/bge-base-en-v1.5 -1.287 0.527 - - -
+BAAI/bge-large-en-v1.5 -2.282 -2.048 - - -
+FacebookAI/roberta-base 1.638 1.989 - 0.23 -0.245
+google-bert/bert-base-uncased 0.585 -1.521 - 0.483 -0.111
+google-bert/bert-large-uncased 1.053 2.75 - - -
+google/gemma-2-2b -5.968 -0.995 - - -
+google/gemma-2b -1.697 0 - - -
+google/gemma-7b -13.517 -2.867 - - -
+GritLM/GritLM-7B -2.458 1.229 - - -
+openai-community/gpt2 -5.5 -4.389 -3.133 - -
+Salesforce/SFR-Embedding-2_R 1.17 1.58 - - -
+state-spaces/mamba-1.4b-hf -5.734 1.989 - - -
+state-spaces/mamba-2.8b-hf -4.096 -1.814 - - -
+Table 9: Downstream testing: $\Delta$ of the Accuracy Score(out of 100) between downstream testing on SST Dataset using the Biased embedding and the Debiased embeddings using various debiasing methods
+Bias Type Model Name $\Delta$ Soft-Debias $\Delta$ DSD $\Delta$ Self-Debias $\Delta$ INLP $\Delta$ SentDebias
+gender Alibaba-NLP/gte-Qwen2-7B-instruct -0.041 -0.04 - - -
+BAAI/bge-base-en-v1.5 -0.06 -0.083 - - -
+BAAI/bge-large-en-v1.5 0.048 -0.033 - - -
+FacebookAI/roberta-base 0.05 -0.029 - 0.029 -0.024
+google-bert/bert-base-uncased -0.055 0.024 - -0.014 -0.024
+google-bert/bert-large-uncased 0.206 0.157 - - -
+google/gemma-2-2b -0.185 -0.149 - - -
+google/gemma-2b -0.075 0.071 - - -
+google/gemma-7b 0.008 0.116 - - -
+GritLM/GritLM-7B -0.069 -0.092 - - -
+openai-community/gpt2-xl -0.07 0.099 0.056 0.068 0.04
+Salesforce/SFR-Embedding-2_R -0.005 -0.074 - - -
+state-spaces/mamba-1.4b-hf -0.027 0.021 - - -
+state-spaces/mamba-2.8b-hf 0.051 -0.011 - - -
+race Alibaba-NLP/gte-Qwen2-7B-instruct 0.016 0.212 - - -
+BAAI/bge-base-en-v1.5 0.068 0.197 - - -
+BAAI/bge-large-en-v1.5 0.04 0.173 - - -
+FacebookAI/roberta-base -0.015 0.035 - 0.011 -0.022
+google-bert/bert-base-uncased 0.022 0.197 - -0.044 0.051
+google-bert/bert-large-uncased -0.005 0.018 - - -
+google/gemma-2-2b 0.043 0.034 - - -
+google/gemma-2b -0.064 -0.042 - - -
+google/gemma-7b 0.078 0.026 - - -
+GritLM/GritLM-7B 0.021 0.143 - - -
+openai-community/gpt2 -0.059 0.069 -0.126 -0.128 0
+openai-community/gpt2-xl -0.114 -0.008 -0.117 -0.114 -0.208
+Salesforce/SFR-Embedding-2_R 0.076 0.018 - - -
+state-spaces/mamba-1.4b-hf -0.026 -0.045 - - -
+state-spaces/mamba-2.8b-hf 0.007 0.08 - - -
+religion Alibaba-NLP/gte-Qwen2-7B-instruct 0.057 0.016 - - -
+BAAI/bge-base-en-v1.5 0.015 0.072 - - -
+BAAI/bge-large-en-v1.5 -0.107 -0.053 - - -
+FacebookAI/roberta-base -0.074 -0.065 - -0.03 0.016
+google-bert/bert-base-uncased -0.101 -0.118 - 0.048 -0.04
+google-bert/bert-large-uncased 0.013 0.154 - - -
+google/gemma-2-2b -0.091 -0.029 - - -
+google/gemma-2b 0.033 0.024 - - -
+google/gemma-7b 0.003 -0.008 - - -
+GritLM/GritLM-7B -0.061 -0.026 - - -
+openai-community/gpt2 -0.095 -0.034 0.092 0.059 0.052
+Salesforce/SFR-Embedding-2_R -0.06 -0.015 - - -
+state-spaces/mamba-1.4b-hf 0.03 -0.005 - - -
+state-spaces/mamba-2.8b-hf 0.07 0.033 - - -
+Table 10: Downstream testing: $\Delta$ of the Pearson Correlation Score between downstream testing on STS-B Dataset using the Biased embedding and the Debiased embeddings using various debiasing methods
+Bias Type Model Name $\Delta$ Soft-Debias $\Delta$ DSD $\Delta$ Self-Debias $\Delta$ INLP $\Delta$ Sent-Debias
+gender Alibaba-NLP/gte-Qwen2-7B-instruct 0.151 0.154 - - -
+BAAI/bge-base-en-v1.5 0.168 0.156 - - -
+BAAI/bge-large-en-v1.5 0.156 0.16 - - -
+FacebookAI/roberta-base 0.014 0.152 - -0.011 0.002
+google-bert/bert-base-uncased -0.151 -0.001 - -0.011 0.002
+google-bert/bert-large-uncased 0.154 -0.007 - - -
+google/gemma-2-2b -0.166 -0.157 - - -
+google/gemma-2b 0.003 -0.145 - - -
+google/gemma-7b 0.005 -0.007 - - -
+GritLM/GritLM-7B -0.155 0.007 - - -
+openai-community/gpt2-xl 0.163 0.166 0.005 0.001 -0.006
+Salesforce/SFR-Embedding-2_R 0.146 0.144 - - -
+state-spaces/mamba-1.4b-hf 0.157 0.005 - - -
+state-spaces/mamba-2.8b-hf 0.077 0.078 - - -
+race Alibaba-NLP/gte-Qwen2-7B-instruct 0.003 -0.139 - - -
+BAAI/bge-base-en-v1.5 0.156 0.167 - - -
+BAAI/bge-large-en-v1.5 -0.001 0.158 - - -
+FacebookAI/roberta-base -0.014 -0.021 - -0.008 0.01
+google-bert/bert-base-uncased 0.011 0.015 - -0.003 0.014
+google-bert/bert-large-uncased -0.175 -0.42 - - -
+google/gemma-2-2b 0.171 -0.239 - - -
+google/gemma-2b 0.174 0.161 - - -
+google/gemma-7b 0.238 0.245 - - -
+GritLM/GritLM-7B -0.006 -0.169 - - -
+openai-community/gpt2 -0.133 0.012 -0.002 -0.002 0.005
+openai-community/gpt2-xl -0.417 -0.018 0.005 0.001 -0.006
+Salesforce/SFR-Embedding-2_R 0 0.813 - - -
+state-spaces/mamba-1.4b-hf 0.404 0.404 - - -
+state-spaces/mamba-2.8b-hf -0.001 -0.164 - - -
+religion Alibaba-NLP/gte-Qwen2-7B-instruct 0.166 0.166 - - -
+BAAI/bge-base-en-v1.5 -0.172 0 - - -
+BAAI/bge-large-en-v1.5 -0.009 0 - - -
+FacebookAI/roberta-base -0.159 -0.006 - -0.003 0.014
+google-bert/bert-base-uncased -0.162 -0.146 - -0.008 0.005
+google-bert/bert-large-uncased 0.001 -0.007 - - -
+google/gemma-2-2b 0.002 -0.159 - - -
+google/gemma-2b 0.009 -0.153 - - -
+google/gemma-7b -0.14 0.015 - - -
+GritLM/GritLM-7B -0.006 -0.014 - - -
+openai-community/gpt2 0.008 0.018 -0.002 -0.002 0.005
+Salesforce/SFR-Embedding-2_R 0.241 0.235 - - -
+state-spaces/mamba-1.4b-hf 0.012 -0.144 - - -
+state-spaces/mamba-2.8b-hf -0.138 0 - - -
+Table 11: Downstream testing: $\Delta$ of the F1 Score between downstream testing on MRPC Dataset using the Biased embedding and the Debiased embeddings using various debiasing methods
+Bias Type Model Name $\Delta$ Soft-Debais $\Delta$ DSD $\Delta$ Self-Debias $\Delta$ INLP $\Delta$ Sent-Debias
+gender Alibaba-NLP/gte-Qwen2-7B-instruct -0.003 -0.012
+BAAI/bge-base-en-v1.5 -0.003 0.005
+BAAI/bge-large-en-v1.5 -0.001 -0.008
+FacebookAI/roberta-base -0.018 0.001 0.001 0.004
+google-bert/bert-base-uncased -0.009 -0.011 -0.008 -0.001
+google-bert/bert-large-uncased -0.003 0
+google/gemma-2-2b -0.002 0.005
+google/gemma-2b -0.002 0.004
+google/gemma-7b 0.004 -0.015
+GritLM/GritLM-7B -0.012 -0.01
+openai-community/gpt2-xl -0.005 -0.004 -0.004 -0.003 0.003
+Salesforce/SFR-Embedding-2_R 0.009 0.009
+state-spaces/mamba-1.4b-hf -0.004 0.012
+state-spaces/mamba-2.8b-hf 0 0.002
+race Alibaba-NLP/gte-Qwen2-7B-instruct 0 0.001
+BAAI/bge-base-en-v1.5 -0.008 0.009
+BAAI/bge-large-en-v1.5 0.003 0.012
+FacebookAI/roberta-base 0.001 -0.003 -0.008 -0.001
+google-bert/bert-base-uncased -0.001 0.001 -0.008 -0.001
+google-bert/bert-large-uncased -0.014 0.001
+google/gemma-2-2b -0.002 0.004
+google/gemma-2b -0.018 -0.021
+google/gemma-7b 0.014 -0.003
+GritLM/GritLM-7B -0.006 -0.005
+openai-community/gpt2 -0.004 0.005 0.004 0.017 0.004
+openai-community/gpt2-xl -0.001 0.005 -0.004 -0.003 0.003
+Salesforce/SFR-Embedding-2_R 0.013 0.012
+state-spaces/mamba-1.4b-hf -0.01 0.009
+state-spaces/mamba-2.8b-hf -0.005 0.001
+religion Alibaba-NLP/gte-Qwen2-7B-instruct -0.011 0.002
+BAAI/bge-base-en-v1.5 -0.002 -0.012
+BAAI/bge-large-en-v1.5 -0.007 0
+FacebookAI/roberta-base 0.006 -0.006 -0.008 -0.001
+google-bert/bert-base-uncased 0 -0.001 -0.005 0.009
+google-bert/bert-large-uncased 0.001 0.008
+google/gemma-2-2b -0.013 -0.012
+google/gemma-2b -0.012 -0.001
+google/gemma-7b 0.003 -0.005
+GritLM/GritLM-7B 0 0.008
+openai-community/gpt2 0.007 0.011 0.004 0.017 0.004
+Salesforce/SFR-Embedding-2_R 0.021 0.011
+state-spaces/mamba-1.4b-hf -0.011 0.001
+state-spaces/mamba-2.8b-hf -0.006 0.002
+Table 12: Downstream testing: $\Delta$ of the F1 Score between downstream testing on MNLI Dataset using the Biased embedding and the Debiased embeddings using various debiasing methods
+Bias Type Model Name $\Delta$ Soft-Debias $\Delta$ DSD $\Delta$ Self-Debias $\Delta$ INLP $\Delta$ Sent-Debias
+gender Alibaba-NLP/gte-Qwen2-7B-instruct 0 0.452
+BAAI/bge-base-en-v1.5 -0.002 0.203
+BAAI/bge-large-en-v1.5 0.005 0.008
+FacebookAI/roberta-base 0.143 0.113 0.016 -0.005
+google-bert/bert-base-uncased -0.099 -0.101 0.033 0.201
+google-bert/bert-large-uncased -0.118 -0.115
+google/gemma-2-2b 0.006 -0.008
+google/gemma-2b -0.092 -0.417
+google/gemma-7b 0.037 -0.002
+GritLM/GritLM-7B 0.325 0.411
+openai-community/gpt2-xl 0.075 -0.333
+Salesforce/SFR-Embedding-2_R 0 0.579
+state-spaces/mamba-1.4b-hf 0.012 0.011
+state-spaces/mamba-2.8b-hf -0.042 -0.029
+race Alibaba-NLP/gte-Qwen2-7B-instruct -0.005 -0.008
+BAAI/bge-base-en-v1.5 0.043 0.015
+BAAI/bge-large-en-v1.5 -0.022 0.089
+FacebookAI/roberta-base 0.002 0.011 0.011 0.005
+google-bert/bert-base-uncased 0.03 0.199 0.018 0.065
+google-bert/bert-large-uncased 0.138 -0.02
+google/gemma-2-2b 0.008 0.142
+google/gemma-2b -0.013 0
+google/gemma-7b 0.02 0.006
+GritLM/GritLM-7B 0.01 0.026
+openai-community/gpt2 -0.088 0.008 0.01
+openai-community/gpt2-xl 0.118 0.019 -0.046
+Salesforce/SFR-Embedding-2_R -0.644 -0.644
+state-spaces/mamba-1.4b-hf -0.005 0.107
+state-spaces/mamba-2.8b-hf -0.013 -0.011
+religion Alibaba-NLP/gte-Qwen2-7B-instruct -0.009 0.066
+BAAI/bge-base-en-v1.5 0.003 0.001
+BAAI/bge-large-en-v1.5 0.021 0.015
+FacebookAI/roberta-base 0.021 0.005 -0.09 -0.134
+google-bert/bert-base-uncased 0.314 0.513 0.004 0.005
+google-bert/bert-large-uncased -0.029 0.065
+google/gemma-2-2b 0.019 0.052
+google/gemma-2b 0.012 0.175
+google/gemma-7b -0.15 -0.178
+GritLM/GritLM-7B -0.009 -0.018
+openai-community/gpt2 0.322 0.338 -0.014
+Salesforce/SFR-Embedding-2_R -0.073 0
+state-spaces/mamba-1.4b-hf 0.007 -0.014
+state-spaces/mamba-2.8b-hf 0.011 0.13
+Table 13: Downstream testing: $\Delta$ of the F1 Score between downstream testing on RTE Dataset using the Biased embedding and the Debiased embeddings using various debiasing methods
+Bias Type Model Name $\Delta$ Soft-Debias $\Delta$ DSD $\Delta$ Self-Debias $\Delta$ INLP $\Delta$ Sent-Debias
+gender Alibaba-NLP/gte-Qwen2-7B-instruct -0.011 0.011
+BAAI/bge-base-en-v1.5 -0.011 0.014
+BAAI/bge-large-en-v1.5 0.033 -0.09
+FacebookAI/roberta-base 0.011 0.021 0.078 -0.339
+google-bert/bert-base-uncased 0 -0.057 0.01 -0.014
+google-bert/bert-large-uncased 0.05 0.064
+google/gemma-2-2b -0.004 -0.011
+google/gemma-2b -0.039 0
+google/gemma-7b 0.03 0.041
+GritLM/GritLM-7B -0.011 0.007
+openai-community/gpt2-xl -0.056 -0.024 0.081 -0.011 0.044
+Salesforce/SFR-Embedding-2_R 0.019 0.004
+state-spaces/mamba-1.4b-hf 0.014 0.017
+state-spaces/mamba-2.8b-hf -0.011 0.007
+race Alibaba-NLP/gte-Qwen2-7B-instruct 0.01 -0.003
+BAAI/bge-base-en-v1.5 -0.018 -0.029
+BAAI/bge-large-en-v1.5 -0.017 -0.131
+FacebookAI/roberta-base 0 0.019 -0.005 0.017
+google-bert/bert-base-uncased -0.051 -0.047 -0.018 -0.263
+google-bert/bert-large-uncased -0.003 -0.028
+google/gemma-2-2b 0.018 0.007
+google/gemma-2b -0.014 0.01
+google/gemma-7b 0.01 0
+GritLM/GritLM-7B -0.018 0
+openai-community/gpt2 0.023 0.015 0.118 -0.221 0.107
+openai-community/gpt2-xl 0.004 -0.019 0.014 0.034 -0.091
+Salesforce/SFR-Embedding-2_R 0.056 -0.024
+state-spaces/mamba-1.4b-hf 0.037 0.015
+state-spaces/mamba-2.8b-hf -0.008 0.037
+religion Alibaba-NLP/gte-Qwen2-7B-instruct -0.035 -0.028
+BAAI/bge-base-en-v1.5 -0.014 0.026
+BAAI/bge-large-en-v1.5 -0.022 0.014
+FacebookAI/roberta-base 0.007 -0.062 -0.059 0.016
+google-bert/bert-base-uncased 0 0.01 -0.018 -0.029
+google-bert/bert-large-uncased -0.007 0
+google/gemma-2-2b 0 -0.021
+google/gemma-2b -0.033 0.006
+google/gemma-7b -0.011 0.004
+GritLM/GritLM-7B -0.037 -0.007
+openai-community/gpt2 -0.011 -0.026 -0.007 0.18 0.003
+Salesforce/SFR-Embedding-2_R 0.035 0.009
+state-spaces/mamba-1.4b-hf -0.018 0.011
+state-spaces/mamba-2.8b-hf 0.032 0.011
+Table 14: Downstream testing: $\Delta$ of the F1 Score between downstream testing on WNLI Dataset using the Biased embedding and the Debiased embeddings using various Debiasing Methods
+Table 15: QNLI
+Bias Type Model Name $\Delta$ Soft-Debias $\Delta$ DSD
+gender Alibaba-NLP/gte-Qwen2-7B-instruct -0.21 0.163
+BAAI/bge-base-en-v1.5 -0.036 -0.077
+BAAI/bge-large-en-v1.5 -0.019 -0.11
+FacebookAI/roberta-base -0.099 -0.02
+google-bert/bert-base-uncased -0.018 -0.109
+google-bert/bert-large-uncased 0.006 -0.025
+google/gemma-2-2b -0.321 0.021
+google/gemma-2b -0.256 0.091
+google/gemma-7b -0.153 0.195
+GritLM/GritLM-7B -0.191 0.014
+meta-llama/Meta-Llama-3-8B -0.158 0.246
+openai-community/gpt2-xl -0.015 -0.011
+Salesforce/SFR-Embedding-2_R -0.12 0.17
+state-spaces/mamba-1.4b-hf -0.32 0.009
+state-spaces/mamba-2.8b-hf -0.276 0.067
+race Alibaba-NLP/gte-Qwen2-7B-instruct -0.199 0.132
+BAAI/bge-base-en-v1.5 -0.027 -0.016
+BAAI/bge-large-en-v1.5 0.003 -0.012
+FacebookAI/roberta-base -0.059 -0.04
+google-bert/bert-base-uncased -0.023 -0.08
+google-bert/bert-large-uncased -0.05 -0.106
+google/gemma-2-2b -0.314 0.038
+google/gemma-2b -0.29 0.077
+google/gemma-7b -0.141 0.163
+GritLM/GritLM-7B -0.173 0.191
+meta-llama/Meta-Llama-3-8B -0.146 0.216
+openai-community/gpt2 -0.272 -0.018
+openai-community/gpt2-xl 0.012 -0.013
+Salesforce/SFR-Embedding-2_R -0.16 0.198
+state-spaces/mamba-1.4b-hf -0.351 -0.057
+state-spaces/mamba-2.8b-hf -0.282 -0.038
+religion Alibaba-NLP/gte-Qwen2-7B-instruct -0.201 0.14
+BAAI/bge-base-en-v1.5 -0.027 -0.015
+BAAI/bge-large-en-v1.5 0.023 0.015
+FacebookAI/roberta-base -0.059 -0.023
+google-bert/bert-base-uncased -0.024 -0.044
+google-bert/bert-large-uncased -0.009 -0.051
+google/gemma-2-2b -0.314 0.042
+google/gemma-2b -0.286 0.039
+google/gemma-7b -0.187 0.166
+GritLM/GritLM-7B -0.159 0.19
+meta-llama/Meta-Llama-3-8B -0.154 0.177
+openai-community/gpt2 -0.274 -0.004
+Salesforce/SFR-Embedding-2_R -0.134 0.23
+state-spaces/mamba-1.4b-hf -0.3 -0.016
+state-spaces/mamba-2.8b-hf -0.245 0.066
+Table 16: Downstream testing: $\Delta$ of the F1 Score between downstream testing on WNLI Dataset using the Biased embedding and the Debiased embeddings using Soft-Debiasing and DeepSoftDebias
 Figures 9 and 10 present an illustration of the results of various downstream tasks and their performance evaluation.
@@
 We can see the varying performance of the 3 different sizes according to the embedding dimension of the LM it is used with.
-Model Variant Bias Direction Biased Text Class. Acc. Debiased Text Class. Acc. Biased NER Macro F1 Avg. Debiased NER Macro F1 Avg.
-Mistral e5-mistral-7b-instruct Gender 0.809 0.810 0.385 0.433
-Zephyr Zephyr-7B-beta 0.806 0.810 0.368 0.429
-Salesforce SFR-Embedding-Mistral 0.807 0.821 0.373 0.428
-Yi Yi-6B 0.748 0.782 0.378 0.423
-Alpaca Alpaca-7B 0.745 0.729 0.349 0.391
-BAAI bge-large-en-v1.5 0.831 0.822 0.456 0.421
-BAAI bge-base-en-v1.5 0.831 0.822 0.453 0.392
-Mistral e5-mistral-7b-instruct Race 0.811 0.817 0.362 0.435
-Zephyr Zephyr-7B-beta 0.806 0.814 0.393 0.443
-Salesforce SFR-Embedding-Mistral 0.810 0.818 0.382 0.413
-Yi Yi-6B 0.748 0.782 0.378 0.423
-Alpaca Alpaca-7B 0.751 0.729 0.358 0.373
-BAAI bge-large-en-v1.5 0.832 0.821 0.473 0.424
-BAAI bge-base-en-v1.5 0.830 0.813 0.457 0.433
-Mistral e5-mistral-7b-instruct Religion 0.808 0.815 0.385 0.437
-Zephyr Zephyr-7B-beta 0.805 0.812 0.389 0.444
-Salesforce SFR-Embedding-Mistral 0.808 0.817 0.389 0.440
-Yi Yi-6B 0.750 0.769 0.384 0.439
-Alpaca Alpaca-7B 0.751 0.726 0.364 0.389
-BAAI bge-large-en-v1.5 0.830 0.816 0.457 0.419
-BAAI bge-base-en-v1.5 0.830 0.817 0.459 0.421
-Table 6: Downstream testing results with embeddings debiased using DeepSoftDebias.
-The first two columns represent results for downstream performance on sentiment analysis.
-The second two columns represent results for downstream performance on NER.
-The best performance is highlighted in bold.
 Model Name Topic STS-B $\uparrow$ Baseline Debiased PCC STS-B $\uparrow$ DeepSoftDebias Debiased PCC SST Biased Acc. SST Baseline Debiased Acc. SST DeepSoftDebias Debiased Acc.
@@
 Gemma, released by Google, further demonstrates the impact of training data size, with both 2B and 7B variants drawing from an extensive pool of up to 6 trillion tokens.
-Appendix F Ablation Experiments
+Appendix F Neural Network Schematic
+Figure 12: A schematic of the Embedding Transforming NN with 2 residual blocks
+Appendix G Hyperparameters for the different LLMs tested
+Model Name Embedding Dim Num Res Blocks LR Batch Size Num Epochs
+Alibaba-NLP/gte-Qwen2-7B-instruct 3584 3 1.00e-5 1024 250
+BAAI/bge-base-en-v1.5 768 1 5.00e-5 2048 100
+BAAI/bge-large-en-v1.5 1024 1 5.00e-5 2048 100
+FacebookAI/roberta-base 768 1 5.00e-5 2048 100
+google-bert/bert-base-uncased 768 1 5.00e-5 2048 100
+google-bert/bert-large-uncased 1024 1 5.00e-5 2048 100
+google/gemma-2-2b 2304 2 5.00e-5 1024 200
+google/gemma-2b 2048 2 5.00e-5 1024 200
+google/gemma-7b 3072 2 5.00e-5 1024 250
+GritLM/GritLM-7B 4096 3 5.00e-5 1024 300
+openai-community/gpt2-xl 1600 1 5.00e-5 2048 150
+Salesforce/SFR-Embedding-2_R 4096 3 1.00e-5 1024 300
+state-spaces/mamba-1.4b-hf 2048 2 5.00e-5 1024 200
+state-spaces/mamba-2.8b-hf 2560 2 5.00e-5 1024 250
+Table 18: Table of the different hyperparameters used with the different LLMs
+Appendix H Ablation Experiments
 In our study, we conduct ablation experiments to assess the effectiveness of various debiasing techniques in the realm of natural language processing.